### v1

# H-STAR: LLM-driven Hybrid SQL-Text Adaptive Reasoning on Tables

Abstract
Tabular reasoning involves interpreting unstructured queries against structured tables, requiring a synthesis of textual understanding and symbolic reasoning. Existing methods rely on either of the approaches and are constrained by their respective limitations. Textual reasoning excels in semantic interpretation unlike symbolic reasoning (SQL logic), but falls short in mathematical reasoning where SQL excels. In this paper, we introduce a novel algorithm H-STAR, comprising table extraction and adaptive reasoning, integrating both symbolic and semantic (text-based) approaches. To enhance evidence extraction, H-STAR employs a multi-view approach, incorporating step-by-step row and column retrieval. It also adapts reasoning strategies based on question types, utilizing symbolic reasoning for quantitative and logical tasks, and semantic reasoning for direct lookup and complex lexical queries. Our extensive experiments demonstrate that H-STAR significantly outperforms state-of-the-art methods across three tabular question-answering (QA) and fact-verification datasets, underscoring its effectiveness and efficiency. [^1] The codes and data are available at: https://github.com/nikhilsab/H-STAR
1 Introduction
Tabular data is one of the most widely used formats for storing structured information in real-world applications. Table-based reasoning presents an inherently challenging problem, requiring logical, mathematical, and textual reasoning over unstructured queries and structured tables Ye et al. (2023). Thus, understanding tabular data has become a significant area of research in natural language processing. Tabular reasoning tasks (illustrated in Figure 1), such as table-based question answering Pasupat and Liang (2015); Nan et al. (2022) and table-based fact verification Chen et al. (2019), have been extensively explored in the past.
Figure 1: An illustration of different tabular reasoning tasks (a) Fact-verification, (b) Short-form QA, and (c) Long-form QA. For each task, a question Q is paired with its answer A, which varies by task. Evidence shows the relevant columns and rows needed to answer the question.
Figure 2: An illustration highlighting the complexity of table reasoning and the need for an integrated approach: (a) Original table, (b) Symbolic reasoning misinterprets the question and returns a value instead of a yes/no response, and (c) Text-based approach fails to solve a math question correctly, leading to an incorrect answer.
Large Language Model (LLM) based prompting methods involving symbolic approaches such as Program of Thought Chen et al. (2023) and textual methods like Chain of Thought (CoT) Wei et al. (2022); Brown et al. (2020) have been used for table reasoning. As shown in Figure 2, these methods often fall short due to the complexities imposed by the intricate mix of numerical, temporal, and textual data, coupled with complex table structures Shi et al. (2023); Liu et al. (2023); Sui et al. (2024). Recent techniques Cheng et al. (2022); Ye et al. (2023); Wang et al. (2023); Nahid and Rafiei (2024) rely on either textual or symbolic reasoning. Textual reasoning excels in natural language understanding but often misinterprets table structures and struggles with quantitative reasoning. Conversely, SQL-based approaches are strong in quantitative problem-solving but perform poorly on noisy or unstructured inputs Liu et al. (2023). Integrating both methods can leverage their complementary strengths, thereby addressing their respective weaknesses. Therefore, it is crucial to explore the question: Can we integrate both symbolic and textual approaches into a hybrid method to enhance tabular reasoning?
In this paper, we introduce H-STAR which integrates symbolic and textual reasoning abilities of LLMs, achieving the best of both worlds. H-STAR decomposes the complex task of table reasoning task into two sub-tasks: a) Table Extraction, where it extracts a question-specific table, and b) Adaptive Reasoning, where it combines SQL-based and text-based reasoning. In the table extraction phase, we employ a step-by-step ‘multi-view’ chain, first identifying relevant columns using both the original and transposed tables. The filtered table is then used for row extraction, contrary to other approaches, which use the original table for both row and column selection in a single view Ye et al. (2023); Nahid and Rafiei (2024).
Focusing on relevant table cells helps LLMs by providing the right context, thereby reducing hallucinations in reasoning. We utilize both SQL-based and text-based techniques in H-STAR’s table extraction phase. In the adaptive reasoning phase, we guide the LLM through a process of step-by-step ‘adaptive’ reasoning, where it chooses between symbolic (SQL) and semantic (textual) methods. H-STAR tailors its reasoning strategies based on the nature of the questions, employing symbolic reasoning for quantitative, mathematical, and logical tasks, while opting for semantic reasoning for direct lookup, common-sense, and challenging lexical queries.
We evaluate the efficacy of H-STAR across three tabular benchmarks involving table QA and table fact verification. Our experiments show that using the combined approach of H-STAR, aligned with multi-view extraction and adaptive reasoning outperforms all prior state-of-the-art approaches. Our results demonstrate that H-STAR is robust and generalizable across LLMs. Through ablations, error analysis, and performance evaluations, we illustrate the effectiveness of H-STAR. Our main contributions are:
(a) We introduce H-STAR, a two-step method unifying table extraction with adaptive reasoning to enhance performance on tabular reasoning. It is a hybrid approach that effectively integrates symbolic (SQL logic) and semantic (textual) methods.
(b) Through extensive experiments, we demonstrate that H-STAR’s hybrid approach outperforms individually implemented symbolic and semantic methods, particularly excelling on longer tables.
(c) Our experiments reveal H-STAR’s efficiency in table extraction using its precise ‘multi-view’ mechanism. Our analysis also highlights H-STAR’s superior reasoning capabilities, emphasizing the necessity of every step.
2 H-STAR Approach
[…]
2.1 Table Extraction
Two-dimensional tables consist of columns (attributes) and rows (values). Table extraction involves a two-step reasoning chain: (1) column extraction and (2) row extraction (see Figure 3). As illustrated in Algorithm 1, for a given table T, table extraction returns a table Tcr (Steps 1-8) based on the input query Q.
Figure 3: An overview of our proposed method, H-STAR, consisting of a combination of code generation and text-based verification. Given a complex table and its question, H-STAR answers using (a) Table Extraction: extracts the question-specific table from the original by first selecting the columns followed by rows. (b) Adaptive Reasoning: when the question has any mathematical component, it generates an additional table using SQL used in the textual reasoning step.
Column Extraction.
As shown in Figure 3, H-STAR uses a ‘multi-view’ technique for column extraction, working with both the original table T (Step 1) and its transposed form TT (Step 2). This ensures that the LLM semantically understands and extracts correct attributes regardless of the table structure. Given an input table T and the question Q (‘How long did it take the New York Americans to win the National Cup after 1936?’), the LLM employs colsql to generate a SQL query to extract columns ‘year’ (C1). This is followed by the text-based column verification step coltext on TT returning a new set of columns C2 (‘year’, ‘national cup’). Table T is filtered for columns ‘year’, ‘national cup’ (C) derived from Steps 1 and 2 to obtain a column-filtered table $\textsc{T}_{\textsc{c}}$ (Step 4).
Algorithm 1 H-STAR
0:  (T, Q) $\rightarrow$ Table - Question pair
0:  $\hat{A}$rightarrow Predicted AnswerTable ExtractionStep 1: Column Extraction
1:   C1$\leftarrow$ colsql(T, Q) $\triangleright$ SQL extraction
2:   C2 $\leftarrow$ coltext(TT, Q, C1) $\triangleright$ Text $\triangleright$ $\textsc{T}^{\textsc{t}}$ refers to the transposed table
[…]
7:   ${}^{\prime}\leftarrow$ R1 $\cup$ R2
8:   TCR $\leftarrow$ TC.filter(R) $\triangleright$ Filter table $\textsc{T}_{\textsc{c}}$ with rows $\textsc{R}^{\prime}$ to get $\textsc{T}_{\textsc{cr}}$daptive Reasoning
9:  if math(Q) ==absent$=$ = True then
10:     Ev $\leftarrow$ fsql(Tcr, Q)
11:     $\hat{A}\leftarrow$ ftext(Tcr, Q, Ev)$\triangleright$ Use SQL generated Ev with Tcr to generate $\hat{A}$.
12:  else
13:     $\hat{A}\leftarrow$ ftext(Tcr, Q) $\triangleright$ Use only the table $\textsc{T}_{\textsc{cr}}$ to generate $\hat{A}$
14:  end if
15:  return  $\hat{A}$
Row Extraction.
After obtaining the filtered table $\textsc{T}_{\textsc{c}}$ with relevant columns, H-STAR proceeds to the row extraction phase culminating in the query-specific table Tcr (Steps 5-8). The filtered table $\textsc{T}_{\textsc{c}}$ has fewer tokens than T, enabling more efficient row extraction. The SQL query generated by rowsql on the column-filtered table $\textsc{T}_{\textsc{c}}$ extracts the relevant rows R1 (‘row 18’). While it returns the appropriate row, it is not sufficient to answer the question. To overcome these limitations, we use text-based row verification rowtext to obtain the missing rows, ultimately retrieving ‘row 1’ and ‘row 18’ ($\textsc{R}_{2}$). Combining $\textsc{R}_{1}$ and $\textsc{R}_{2}$ gives the relevant rows $\textsc{R}^{\prime}$ (‘row 1’, ‘row 18’), resulting in the final table $\textsc{T}_{\textsc{cr}}$.
2.2 Adaptive Reasoning
To leverage the strengths of textual reasoning while overcoming its quantitative limitations, we introduce an adaptive reasoning framework. Assessing the nature of the question, the LLM incorporates symbolic reasoning only when required. For instance, in Figure 3, when faced with a question like "How long did it take the New York Americans to win the National Cup after 1936?", LLM requires a count as an answer, and our pipeline prioritizes symbolic reasoning (fsql) through SQL-generated code. This code is executed on a SQL engine, enabling precise answers from the table $\textsc{T}_{\textsc{cr}}$. The output from this SQL-based evaluation Ev serves as additional evidence, enhancing the textual reasoning process (ftext). By explicitly integrating this quantitative approach, our algorithm effectively addresses LLMs’ limitations in handling mathematical operations, thereby improving the accuracy and robustness when tackling questions that demand numerical reasoning.
3 Experiments
Benchmark Datasets.
We evaluate our method on three datasets covering fact verification, short-form, and long-form question-answering tasks using in-context examples. (a) TabFact Chen et al. (2019): A fact verification benchmark utilizing Wikipedia tables. We evaluate the test-small set, containing 2,024 statements and 298 tables, with each statement labeled as Entailed ("True") or Refuted ("False"). (b) WikiTQ Pasupat and Liang (2015): The WikiTableQuestions (WikiTQ) dataset involves question-answering tasks over semi-structured Wikipedia tables. It includes a standard test set with 4,344 table-question pairs. (c) FeTaQA Nan et al. (2022): FeTaQA (Free-Form Table Question Answering) comprises free-form questions requiring synthesis from multiple table sections for responses. It demands advanced reasoning and profound tabular data understanding, evaluated on a test set of 2003 samples.
Evaluation Metrics.
We tailor our evaluation metrics based on the task and dataset characteristics. For fact-verification datasets like TabFact, we evaluate performance using binary classification accuracy. For short-form question-answering datasets such as WikiTQ, we assess accuracy by measuring the exact match between predicted outputs and the gold-standard answers. For more complex tasks in FeTaQA, which involve long-form question answering, we evaluate performance using ROUGE-1, ROUGE-2, and ROUGE-L metrics Lin (2004), comparing predicted outputs against the long-form answers.
LLM Models.
In our research, we use state-of-the-art large language models (LLM) such as PaLM-2 Anil et al. (2023) and GPT-3.5-Turbo Brown et al. (2020); OpenAI (2023) for table reasoning tasks. Our model inputs include in-context examples, the table, and the question for each step of the pipeline. Previous studies, such as Chain-of-Table and TabSQLify, influenced our model choices to ensure a fair comparison. All baselines with PALM-2 are from the propriety PALM-2 model, which is different from the public PALM-2 API (used in H-STAR and TabSQLify).
Baseline Methods.
We compare our method with (a) generic reasoning based on language models, and (b) table-manipulation reasoning based on language models.
(a) Generic Reasoning. These methods direct the LLM to carry out the required downstream task based on the information from the table and the input question. They include End-to-End QA, which offers only the table and the question, and Few-Shot QA, which involves a few examples with the table-question-answer triplet alongside the table and the question. Chain-of-Thought prompts the LLMs to provide a supporting reasoning chain that leads to the answer.
(b) Table Manipulation. These techniques involve several steps, with the initial stage dedicated to automatically pre-processing the table for the reasoning task. BINDER utilizes in-context examples to produce SQL or Python code containing LLM API calls to generate answers. DATER directs the language model to break down tables and questions, applying reasoning to the decomposed tables based on sub-questions. Chain-of-Table employs the table as an intermediate output in its reasoning process, iterating through tasks until the final answer is obtained. TabSQLify uses SQL to trim the table and then reasons based on the reduced table.
3.1 Results and Analysis
Table 1 compares the performance of different methods on the table reasoning tasks, TabFact, and WikiTQ datasets, across several LLMs including GPT-3.5-Turbo and PaLM-2. This comparison involves evaluating against generic reasoning, table manipulation techniques, and our method, H-STAR. Appendix A provides results on the FeTaQA dataset in addition to comparisons with other baseline models on the WikiTQ and TabFact dataset.
GPT-3.5-Turbo PaLM-2
TabFact WikiTQ TabFact WikiTQ
Generic Reasoning
End-to-End QA 70.45 51.84 77.92 60.59
Few-shot QA 71.54 52.56 78.06 60.33
CoT 65.37 53.48 79.05 60.43
Table Manipulation
BINDER 79.17 56.74 76.98 54.88
DATER 78.01 52.90 84.63 61.48
Chain-of-Table* 80.20 59.94 86.61 67.31
TabSQLify 79.50 64.70 79.78 55.78
H-STAR 83.74 68.85 85.92 68.51
Table 1: Comparison of various methods across datasets. The results are reported from the Chain-of-Table paper for fair comparison. * Chain-of-Table uses a proprietary PaLM-2 model that is better than their publicly available one.
Analysis. (a) On the WikiTQ dataset, H-STAR demonstrates superior performance with both GPT-3.5-Turbo and PaLM-2 models, surpassing all baseline methods. Specifically, H-STAR achieves an accuracy of 68.85% on WikiTQ with GPT-3.5-Turbo, marking a significant improvement of 17.01% over the vanilla GPT-3.5-Turbo model and a 4% increase over the previous best method, TabSQLify. (b) On the TabFact dataset, H-STAR outperforms all methods with GPT-3.5-Turbo, attaining an accuracy of 83.74%, which is a 13.29% improvement over the vanilla model and a 4% margin over the previous best, Chain-of-Table. However, with PaLM-2, H-STAR achieves comparable performance on TabFact while outperforming Chain-of-Table by 1.2% on WikiTQ.
Efficiency of Table Extraction.
Stage-1 in H-STAR involves the extraction of the table most relevant to the question. Our table extraction method uses a two-step chain to select the relevant columns followed by the rows. Figure 4 showcases the effectiveness of the technique to extract only the relevant information from the table. Figure 4 compares the number of average table cells for the extracted table for H-STAR, H-STAR without row extraction, and H-STAR without column extraction with other baselines.
Analysis. In particular, H-STAR significantly reduces the average number of processed table cells - from 159 to 18 for the WikiTQ dataset, from 88 to 13 for the TabFact dataset, and from 86 to 10 for the FeTaQA dataset. In contrast, other table manipulation methods, like TabSQLify and DATER yield higher values. Our approach showcases superior effectiveness, notably reducing cell counts across all datasets. This emphasizes how H-STAR effectively reduces irrelevant information, thus focusing LLMs on the right evidence for right reasoning.
Figure 4: Comparison of average table cells in the final table.
Performance on Longer Tables.
Table-CoT Chen (2023) highlights that longer table size, with tables exceeding 30 rows, is a significant cause of erroneous generations. This observation is supported by a decline in LLM performance as the number of tokens in the table increases. We classify LLM performance based on total tokens into three groups - small (< 2000 tokens), medium (2000 to 4000 tokens), and large (> 4000 tokens). We then compare H-STAR with other table manipulation methods, including BINDER, DATER, Chain-of-Table, and TabSQLify.
Method Small Medium Large
BINDER 56.54 25.13 6.41
DATER 62.50 42.34 34.62
Chain-of-Table 68.13 52.25 44.87
TabSQLify 68.15 57.91 52.34
H-STAR 71.19 64.07 62.50
Table 2: Performance of various methods on different table sizes on WikiTQ.
Analysis. The results shown in Table 2 emphasize the challenges LLMs face when reasoning over longer tables. Our findings align with previous research, revealing a significant decrease in LLM performance as table sizes increase. While other table manipulation methods experience a drop in performance with larger tables, H-STAR maintains consistent performance across all table sizes.
Figure 5: Comparison of SQL-based and text-based column extraction methods for answering table queries: (a) SQL-based selection accurately identifies the columns ("1940/41", "description losses"), while text-based selection mistakenly includes "murdered" as a table column. (b) The text-based method selects both required columns ("year" and "national cup"), while the SQL-based approach overlooks "national cup" as a requirement.
Figure 6: Comparison of SQL-based and text-based row extraction methods for answering table queries. (a) SQL-based selection accurately selects the necessary rows ("Murdered"), while text-based selection incorrectly includes additional rows; (b) Text-based method correctly selects relevant rows ("1953/54", "1936/37"), while SQL-based selection misinterprets the query, including only the row for the year "1953/54".
3.2 Efficiency Analysis
In Table 3, we study H-STAR by considering the total number of samples generated by the LLMs. BINDER and DATER utilize self-consistency techniques to refine their results, while the Chain-of-Table approach adopts a more iterative sample creation process. Specifically, BINDER generates 50 Neural-SQL samples, implementing self-consistency, while DATER employs self-consistency at each step, resulting in 100 samples. In contrast, the Chain-of-Table method opts for a more resource-efficient strategy, producing 25 samples across three steps: ‘Dynamic Plan’, ‘Generate Args’, and ‘Query’. TabSQLify generates the fewest samples, with one generation each for the table decomposition and query steps. Using single outputs can result in a decrease in accuracy when the model fails to return an answer. Thus, H-STAR incorporates self-consistency measures by generating two outputs at each stage to ensure output validity. Each step, except Query, involves two generations for both SQL and Text, i.e., a total of four steps for column and row retrieval. These additional steps serve as a precaution against abnormal LLM generations. Overall, H-STAR is efficient and only uses 6-10 sample generations.
Method # samples / step Total # samples
BINDER Neural SQL: 50 50
DATER Decompose Table: 40 100
Generate Cloze: 20
Generate SQL: 20
Query: 20
Chain-of-Table Dynamic Plan $\leq$ 5 $\leq$ 25
Generate Args $\leq$ 19
Query: 1
TabSQLify Table Decompose: 1 2
Query: 1
H-STAR Column Retrieval: 2-4 6-10
Row Retrieval: 2-4
Query: 2
Table 3: Number of generated samples for different methods.
4 Ablation Study
We conduct an ablation study to measure the contribution of each step in the pipeline and assess the impact of integrating symbolic-textual and adaptive reasoning in H-STAR.
4.1 Importance of Each Stage
To evaluate the significance of the two primary stages in H-STAR: a) Table extraction and b) Adaptive reasoning, we conducted an ablation study. This study involved systematically removing one stage at a time. For the first condition, the final adaptive reasoning stage was substituted with a basic chain-of-thought reasoning process, while maintaining the table extraction stage. Conversely, the second condition retained the adaptive reasoning stage but omitted the table extraction stage, specifically by removing the row extraction and column extraction steps individually and then collectively. A detailed assessment of the contribution of each stage to the overall performance of the method is shown in Table 4. It highlights the importance of each step in the pipeline. Removing any step can result in a performance loss except for a marginal increase in the case of row extraction for TabFact.
Method TabFact WikiTQ
H-STAR 85.92 68.51
w/o row extraction 86.17 66.30
w/o column extraction 84.04 67.03
w/o table extraction 83.79 63.58
w/o adaptive reasoning 79.35 61.47
Table 4: Performance on TabFact and WikiTQ datasets.
4.2 Impact of a Hybrid SQL-Text Approach
We perform a quantitative evaluation of the impact of the hybrid approach in H-STAR. Figures 5, 6 depict how using only one of the approaches results in insufficient outputs. Furthermore, it is evident from Table 5 that not employing a hybrid approach (H-STAR) results in a significant performance loss.
Method TabFact WikiTQ
CoT 79.05 60.43
H-STAR ${}_{\text{text}}$ 79.43 61.47
Text-to-SQL 52.05 42.03
H-STAR ${}_{\text{sql}}$ 58.50 46.09
H-STAR 85.93 68.51
Table 5: Performance using only text-based/SQL-based reasoning; H-STAR text: H-STAR with textual reasoning only; H-STAR sql: H-STAR with symbolic reasoning only.
The low accuracy of SQL approaches can be attributed to the incorrect final answers. The text-based reasoning approach achieves reasonably high scores on TabFact and WikiTQ, albeit lower than the hybrid approach. By combining textual and structured reasoning capabilities, the hybrid approach i.e., H-STAR can effectively handle diverse data formats and task requirements, leading to superior performance across various benchmarks.
5 Where do LLMs fail?
This section examines errors in H-STAR, contrasts the failure points with existing approaches, and summarizes our key findings.
5.1 Error Analysis on H-STAR
The disjoint, step-wise nature of H-STAR enables identifying and analyzing failures. Table 6 shows 100 randomly selected incorrect answer instances from the Tabfact and WikiTQ datasets. In this study, ‘Missing Columns’ and ‘Missing Rows’ refer to missing necessary columns and rows, respectively. ‘Incorrect Reasoning’ occurs when H-STAR extracts the correct table, but LLM fails to produce the correct answer. ‘Incorrect Annotations’ include semantically identical answers in different formats, ambiguous questions, and incorrect gold answers. Table 6 shows that for 100 TabFact failures, 2% are missing columns, 9% missing rows, 79% incorrect LLM reasoning, and 10% incorrect annotations. For WikiTQ, 6% are missing columns, 17% missing rows, 54% incorrect reasoning, and 23% incorrect annotations.
Tabfact WikiTQ
Missing Columns 2% 6%
Missing Rows 9% 17%
Incorrect Reasoning 79% 54%
Incorrect Annotation 10% 23%
Table 6: Error analysis on 100 error samples across datasets showcasing the distribution of errors for H-STAR.
Analysis. Fewer errors in column and row extraction indicate that H-STAR efficiently extracts the necessary table. By reducing errors in the table extraction process, we effectively "shift left" addressing problems earlier in the pipeline and enhancing overall performance.
5.2 More Error Analysis: H-STAR vs Others
Table 7 compares errors in H-STAR, with TabSQLify, and BINDER. We use 100 WikiTQ samples where TabSQLify fails. Many errors with TabSQLify (6/100 missing columns, 56/94 missing rows) result from incorrect row extraction. Out of 100, TabSQLify correctly extracts the table for 38 samples, but incorrectly reasons on 29 (76%). BINDER does not extract tables but generates multiple neural SQL queries for table reasoning tasks.
TabSQLify BINDER H-STAR
Missing Columns 6/100 N/A 4/100
Missing Rows 56/94 N/A 16/96
Incorrect Reasoning 29/38 70/100 27/80
Incorrect Annotation 9/9 9/30 10/53
Correct Answers 0/0 21/21 43/43
Table 7: Analysis of error types in 100 samples from WikiTQ where TabSQLify fails. Note: The denominator refers to the remaining samples; we start with 100 samples and subtract the number of error samples to determine the remaining samples.
However, its output shows that it fails to reason correctly in 70 out of 100 cases. In contrast, H-STAR proves more effective in both table extraction and reasoning, achieving correct generations in 43 samples. Additionally, it demonstrates fewer errors in table extraction (4/100 columns; 16/96 rows) and reasoning (27/80, i.e., 34% of samples).
5.3 What did we learn?
Firstly, our evaluation demonstrates that our hybrid approach achieves substantial improvements, surpassing the performance of previous state-of-the-art methods across various table reasoning tasks. Secondly, the quantitative analysis demonstrates that our ‘multi-view’ approach extracts tables with reduced irrelevant data, resulting in more concise tables. Furthermore, the qualitative analysis highlights fewer errors in our table extraction method compared to prior approaches, confirming a decrease in extraneous information.
Thirdly, our analysis indicates consistent performance even with longer tables, emphasizing our method’s effectiveness in accurately extracting relevant information by filtering out noise. The adaptive reasoning component boosts performance, especially for longer tables. Lastly, the ablation study shows that breaking down the task into sub-tasks significantly enhances the overall performance, with each sub-task playing a crucial role in achieving superior results. Moreover, it highlights the constraints of relying solely on either text or SQL approaches, which are effectively addressed by our H-STAR approach. Together, these findings emphasize the substantial advantages of our hybrid and multi-view approach in addressing complex table reasoning tasks.
6 Related Work
Table reasoning tasks require the ability to reason over unstructured queries and structured or semi-structured tables. Traditional approaches like TAPAS Herzig et al. (2020), TAPEX Liu et al. (2021), TABERT Yin et al. (2020), TURL Deng et al. (2022), PASTA Gu et al. (2022) work on pre-training language models jointly on large-scale tabular and text data to reason in an end-to-end manner. Advancements in LLMs allow them to learn from in-context examples, reducing inference costs. Text-to-SQL Rajkumar et al. (2022) and Program-of-Thought Chen et al. (2023) use symbolic methods to solve table-based tasks via Python/SQL programs. Textual-based reasoning techniques like Table-CoT Chen (2023) and Tab-CoT Ziqi and Lu (2023) extend prompting methods such as zero and few-shot CoT for tabular reasoning.
Decomposing problems into smaller, manageable tasks has proven effective in solving complex reasoning challenges Zhou et al. (2022); Khot et al. (2022). Recent techniques in table reasoning follow this approach, either by breaking tasks into fixed sub-tasks Cheng et al. (2022); Ye et al. (2023); Nahid and Rafiei (2024) or by employing iterative methods Jiang et al. (2023); Wang et al. (2023). BINDER Cheng et al. (2022) is an SQL-based approach that modifies SQL statements to include LLM API calls within SQL statements. DATER Ye et al. (2023) and TabSQLify Nahid and Rafiei (2024) are table decomposition methods that use semantic and symbolic reasoning respectively. Chain-of-Table Wang et al. (2023) uses a textual reasoning approach to update tables iteratively. The final table post-update is used for final reasoning. Unlike H-STAR, these techniques depend either on symbolic or semantic reasoning but not both.
7 Conclusion
In this study, we present H-STAR, a novel method that effectively integrates semantic and symbolic approaches, demonstrating superior performance compared to existing methods in tasks involving table reasoning. Our approach involves a two-step LLM-driven process: firstly, employing ‘multi-view’ table extraction to retrieve tables relevant to a query, and then, implementing adaptive reasoning to select the optimal reasoning strategy based on the input query. We address prior bottlenecks with efficient extraction and reasoning, leading to improved overall performance, particularly on longer tables. Our results highlight the need to move beyond relying solely on either technique and demonstrate the effectiveness of an integrated approach that combines the advantages of both methods. Future directions involve testing the adaptability of our methods to semi-structured, complex hierarchical, and relational tables. Enhancing the reasoning process through techniques such as self-consistency and self-verification shows promising potential.
Limitations
Our current work has focused primarily on a subset of table reasoning tasks using datasets sourced from Wikipedia. While this has laid a solid foundation, it limits exploration into diverse reasoning tasks such as table manipulation, text-to-table generation, and table augmentation, which could provide valuable insights and enhance our approach’s capabilities. Additionally, our method’s generalizability is confined to Wikipedia-based datasets, restricting its application to other domains that require specific domain knowledge, which our current approach lacks. Extending our approach to different domains may necessitate integrating domain-specific knowledge to ensure effective reasoning.
Furthermore, our evaluation has been limited to relatively straightforward table structures. Handling more complex data representations such as semi-structured tables, hierarchical tables, and relational databases remains unexplored territory. These structures present unique challenges that our current approach may not effectively address. Finally, our study has focused solely on the English language, potentially limiting applicability to languages with different linguistic complexities that we have not accounted for.
Ethics Statement
We, the authors, affirm that our work adheres to the highest ethical standards in research and publication. We have carefully considered and addressed various ethical issues to ensure the responsible and fair use of computational linguistics methodologies. To facilitate reproducibility, we provide detailed information, including code, datasets (all publicly available and in compliance with their respective ethical standards), and other relevant resources. Our claims align with the experimental results, though some stochasticity is expected with black-box large language models, which we minimize by maintaining a fixed temperature. We provide comprehensive details on annotations, dataset splits, models used, and prompting methods, ensuring our work can be reliably reproduced.
Acknowledgments
We thank the authors of Tabsqlify and BINDER for sharing their code, and Varun Yerram for his help in setting up the implementation process. This research was supported by the Army Research Office under Grant Number W911NF-20-1-0080 and partially funded by ONR Contract N00014-19-1-2620. Lastly, we appreciate the reviewing team’s insightful comments.
Appendix A Additional Results
Tables 8 and 9 show results of other baseline models on WikiTQ and TabFact datasets respectively. Our method, H-STAR, outperformed all other models by substantial margins.
Model Method Accuracy
PLM TaPas 48.8
GraPPa 52.7
LEVER 62.9
ITR 63.4
Codex [^2] OpenAI Codex model is not available publically anymore. GPT-3 CoT 45.7
TableCoT 48.8
DATER 65.9
BINDER 61.9
ReAcTable 65.8
GPT-3.5-Turbo ReAcTable 52.5
TableCoT 52.4
StructGPT 52.2
GPT-3.5-Turbo H-STAR 68.8
Table 8: Comparison of H-STAR with additional methods on WikiTQ dataset.
Model Method Accuracy
PLM Table-BERT 68.1
LogicFactChecker 74.3
SAT 75.5
TaPas 83.9
TAPEX 85.9
SaMoE 86.7
PASTA 90.8
Codex TableCoT 72.6
DATER 85.6
BINDER 85.1
ReAcTable 83.1
GPT-3.5-Turbo ReAcTable 73.1
TableCoT 73.1
GPT-3.5-Turbo H-STAR 83.7
Table 9: Comparison of H-STAR with additional methods on TabFact dataset.
Prompting ROUGE-1 ROUGE-2 ROUGE-L
PaLM-2*
DATER 0.63 0.41 0.53
Chain-of-Table 0.66 0.44 0.56
GPT-3.5-Turbo
Table-CoT 0.62 0.39 0.51
Tabsqlify 0.58 0.35 0.48
H-STAR 0.62 0.39 0.52
Table 10: Performance on FeTaQA dataset. PaLM-2* refers to the Google proprietary PaLM-2 model
Method Fluent Correct Adequate Faithful
T5-large 94.6 54.8 50.4 50.4
Human 95 92.4 95.6 95.6
TableCoT 96 82 75 87
Tabsqlify 97 88 84 93
H-STAR 96.6 87.6 89.6 94
Table 11: Human evaluation results on FeTaQA dataset.
A.1 FeTaQA
Represented in Table 10 are the results for the FeTaQA dataset using the ROUGE scores Lin (2004). We observe an incremental improvement in the scores when compared to other methods on the GPT-3.5-Turbo. The ROUGE scores focus on lexical similarities while ignoring the semantic similarity between predicted and gold outputs. These metrics are known to be insensitive to capturing improvements when using in-context learning since the model does not learn the expected style of the long-form text just from an instruction or a few examples.
Figure 7 illustrates an example where the generated output, despite being correct in answering the question, is penalized by the ROUGE metric. This highlights the limitations of the metric in evaluating the correctness of the generated responses.
Figure 7: An example of FeTaQA representing low ROUGE scores for correct answers.
Human Evaluation
We further evaluate the quality of our generations using human evaluation. We refer to Nan et al. (2022) and evaluate our generations on four factors: 1) Fluency based on the nature and grammar of the output; 2) Correctness measuring the degree of correctness of the prediction; 3) Adequacy to measure if all the necessary information is contained in the output; 4) Faithfulness to check whether the generation is grounded to the final extracted table used for reasoning. We ask five internal human annotators to assign a score of 1-5 for each criterion and report the percentage of samples having values of 4 or 5. Table 11 shows the results by five annotators for our method and compares it with other prior methods. The performance pattern of H-STAR is similar to human performance with comparable fluency and faithfulness while matching the trend of having higher adequacy over correctness. While we marginally outperform other methods, there is still scope for improvement.
Qualitative Analysis
In this section, we perform a qualitative analysis on 100 samples of FeTaQA with TabSQLify Nahid and Rafiei (2024). The study aims to compare the generations of H-STAR with TabSQLify by evaluating the answer quality. The answers are evaluated based on their fluency and adequacy when compared with the gold answers. Table 12 shows that H-STAR has better quality generations in 26% samples, as opposed to only 18% for TabSQLify. ‘Both’ indicates samples where both the algorithms perform on par. This can be attributed to the use of the same LLM. Most of the failures for TabSQLify are derived from an insufficient table, which further highlights the limitations of using a single-view approach for table extraction. Figure 8 showcases such an example where TabSQLify returns ‘information not provided’ because of incorrect of an extracted table. It can be noted that H-STAR generates an answer comparable with the ground truth for the same. Figure 9 visually depicts the outputs from both methods. It illustrates the evaluation process undertaken to compare the quality of the outputs.
Method % of samples
Both 56%
TabSQLify 18%
H-STAR 26%
Table 12: Comparison of H-STAR with TabSQLify on 100 samples.
Figure 8: An example of FeTaQA dataset where TabSQLify fails as a result of insufficient table evidence.
Figure 9: Comparison of the outputs with TabSQLify.
Appendix B Input Table Format
The input table format changes depending on the type of reasoning used. Since H-STAR uses a hybrid approach, the input table format also varies depending on the type of reasoning used. Besides the question and the in-context examples, the table is accompanied by a list of columns and a table caption if available. Providing more context aids in better semantic understanding Singha et al. (2023); Sui et al. (2024).
B.1 Symbolic Reasoning
We adopt the table prompt format from previous SQL-based methods Text-to-SQL Rajkumar et al. (2022); Cheng et al. (2022); Nahid and Rafiei (2024). We include (1) the table schema containing CREATE TABLE followed by the schema, (2) the table header and all the rows separated by tabs, and (3) the list of columns along with the corresponding query. If the prompt exceeds the context limit, we truncate the table rows to fit within the limit. Example of the input table:
CREATE TABLE 2012–13 Exeter City F.C. season(
row_id int,
name text,
league int,
total int)
All rows of the table:
SELECT * FROM w;
row_id name league total
1 danny coles 3 3
4 john o’flynn 11 12
8 jamie cureton 20 20
columns: [’name’, ’league’, ’total’]
B.2 Textual Reasoning
We convert the tables into linear, sequential text. Continuing with the format of previous textual reasoning methods Ye et al. (2023); Chen (2023); Wang et al. (2023), we adopt the PIPE encoding i.e. plain text separated by ‘|’. Furthermore, we append the table with the caption and the list of columns similar to the format in Section B.1. Example of the input table:
table caption: 2012–13 Exeter City F.C. season
col : name | league | total
row 1: danny coles | 3 | 3
row 4: john o’flynn | 11 | 12
row 8: jamie cureton | 20 | 20
columns: [’name’, ’league’, ’total’]
Appendix C Hyperparameters
Table 13 provides the details of PaLM-2 hyper-parameters used for the WikiTQ and TabFact datasets whereas Table 14 showcases the hyperparameter setting for GPT-3.5-Turbo. The hyperparameters ‘samples’ indicate the number of outputs taken for each step of the pipeline whereas the ‘examples’ indicate the number of few-shot demonstrations for each step.
Function temperature top_p output_tokens samples examples
WikiTQ
Colsql 0.4 1.0 512 2 3
Coltext 0.7 1.0 512 2 3
Rowsql 0.4 1.0 512 2 3
Rowtext 0.7 1.0 512 2 2
Reasonsql 0.1 1.0 512 1 3
Reasontext 0.1 1.0 256 1 4
TabFact
Colsql 0.4 1.0 512 2 4
Coltext 0.7 1.0 512 2 3
Rowsql 0.4 1.0 512 2 4
Rowtext 0.7 1.0 512 2 3
Reasonsql 0.1 1.0 512 1 3
Reasontext 0.1 1.0 256 1 5
Table 13: Hyperparameter settings for PaLM-2
Function temperature top_p output_tokens samples examples
WikiTQ
Colsql 0.3 1.0 512 2 3
Coltext 0.4 1.0 512 2 2
Rowsql 0.4 1.0 512 2 3
Rowtext 0.4 1.0 512 2 3
Reasonsql 0.1 1.0 512 1 4
Reasontext 0.1 1.0 256 1 4
TabFact
Colsql 0.2 1.0 512 2 4
Coltext 0.4 1.0 512 2 3
Rowsql 0.4 1.0 512 2 4
Rowtext 0.5 1.0 512 2 3
Reasonsql 0.1 1.0 512 1 4
Reasontext 0.1 1.0 256 1 5
Table 14: Hyperparameter settings for GPT-3.5-Turbo

### v2

@@
 Abstract
-Tabular reasoning involves interpreting unstructured queries against structured tables, requiring a synthesis of textual understanding and symbolic reasoning.
-Existing methods rely on either of the approaches and are constrained by their respective limitations.
-Textual reasoning excels in semantic interpretation unlike symbolic reasoning (SQL logic), but falls short in mathematical reasoning where SQL excels.
-In this paper, we introduce a novel algorithm H-STAR, comprising table extraction and adaptive reasoning, integrating both symbolic and semantic (text-based) approaches.
-To enhance evidence extraction, H-STAR employs a multi-view approach, incorporating step-by-step row and column retrieval.
-It also adapts reasoning strategies based on question types, utilizing symbolic reasoning for quantitative and logical tasks, and semantic reasoning for direct lookup and complex lexical queries.
+Tabular reasoning involves interpreting natural language queries about tabular data, which presents a unique challenge of combining language understanding with structured data analysis.
+Existing methods employ either textual reasoning, which excels in semantic interpretation but struggles with mathematical operations, or symbolic reasoning, which handles computations well but lacks semantic understanding.
+This paper introduces a novel algorithm H-STAR that integrates both symbolic and semantic (textual) approaches in a two-stage process to address these limitations.
+H-STAR employs: (1) step-wise table extraction using ‘multi-view’ column retrieval followed by row extraction, and (2) adaptive reasoning that adapts reasoning strategies based on question types, utilizing semantic reasoning for direct lookup and complex lexical queries while augmenting textual reasoning with symbolic reasoning support for quantitative and logical tasks.
 Our extensive experiments demonstrate that H-STAR significantly outperforms state-of-the-art methods across three tabular question-answering (QA) and fact-verification datasets, underscoring its effectiveness and efficiency.
@@
 (2023).
-Thus, understanding tabular data has become a significant area of research in natural language processing.
+Thus, understanding and inferring tabular data has become a significant area of research in natural language processing.
 Tabular reasoning tasks (illustrated in Figure 1), such as table-based question answering Pasupat and Liang (2015); Nan et al.
@@
 Figure 2: An illustration highlighting the complexity of table reasoning and the need for an integrated approach: (a) Original table, (b) Symbolic reasoning misinterprets the question and returns a value instead of a yes/no response, and (c) Text-based approach fails to solve a math question correctly, leading to an incorrect answer.
-Large Language Model (LLM) based prompting methods involving symbolic approaches such as Program of Thought Chen et al.
+Advancements in Large Language Models (LLMs) have led to better performance across various tasks through carefully crafted prompts.
+In the domain of table reasoning, symbolic approaches such as Program of Thought Chen et al.
 (2023) and textual methods like Chain of Thought (CoT) Wei et al.
 (2022); Brown et al.
-(2020) have been used for table reasoning.
-As shown in Figure 2, these methods often fall short due to the complexities imposed by the intricate mix of numerical, temporal, and textual data, coupled with complex table structures Shi et al.
+(2020) have been explored.
+However, when applied individually (see Figure 2) these methods often struggle due to the complexities imposed by the intricate mix of numerical, temporal, and textual data, coupled with complex table structures Shi et al.
 (2023); Liu et al.
@@
 (2024).
-Recent techniques Cheng et al.
-(2022); Ye et al.
-(2023); Wang et al.
-(2023); Nahid and Rafiei (2024) rely on either textual or symbolic reasoning.
 Textual reasoning excels in natural language understanding but often misinterprets table structures and struggles with quantitative reasoning.
@@
 (2023).
-Integrating both methods can leverage their complementary strengths, thereby addressing their respective weaknesses.
-Therefore, it is crucial to explore the question: Can we integrate both symbolic and textual approaches into a hybrid method to enhance tabular reasoning?
-In this paper, we introduce H-STAR which integrates symbolic and textual reasoning abilities of LLMs, achieving the best of both worlds.
-H-STAR decomposes the complex task of table reasoning task into two sub-tasks: a) Table Extraction, where it extracts a question-specific table, and b) Adaptive Reasoning, where it combines SQL-based and text-based reasoning.
-In the table extraction phase, we employ a step-by-step ‘multi-view’ chain, first identifying relevant columns using both the original and transposed tables.
-The filtered table is then used for row extraction, contrary to other approaches, which use the original table for both row and column selection in a single view Ye et al.
+As a result, recent techniques Cheng et al.
+(2022); Ye et al.
+(2023); Wang et al.
+(2023); Nahid and Rafiei (2024) that rely solely on textual or symbolic reasoning suffer from similar limitations.
+Therefore, it is crucial to explore the question: How can symbolic and textual approaches be integrated into a hybrid method for tabular reasoning that leverages their complementary strengths while mitigating their individual limitations?
+In this paper, we introduce H-STAR which combines symbolic and textual reasoning abilities of LLMs, achieving the best of both worlds.
+H-STAR decomposes the table reasoning task into two sub-tasks: (a) Table Extraction and (b) Adaptive Reasoning.
+In the table extraction phase, we employ a step-by-step ‘multi-view’ chain, first identifying relevant columns using the original table and its transposed form.
+The filtered table is then used for row extraction, contrary to other approaches, which use the original table for both row and column selection Ye et al.
 (2023); Nahid and Rafiei (2024).
 Focusing on relevant table cells helps LLMs by providing the right context, thereby reducing hallucinations in reasoning.
-We utilize both SQL-based and text-based techniques in H-STAR’s table extraction phase.
-In the adaptive reasoning phase, we guide the LLM through a process of step-by-step ‘adaptive’ reasoning, where it chooses between symbolic (SQL) and semantic (textual) methods.
-H-STAR tailors its reasoning strategies based on the nature of the questions, employing symbolic reasoning for quantitative, mathematical, and logical tasks, while opting for semantic reasoning for direct lookup, common-sense, and challenging lexical queries.
-We evaluate the efficacy of H-STAR across three tabular benchmarks involving table QA and table fact verification.
+We employ a combination of SQL-based and text-based techniques in H-STAR’s table extraction phase.
+In the adaptive reasoning phase, we use the language comprehension capabilities of LLMs to guide them with few-shot examples to decide when to support semantic methods with symbolic (SQL) approaches.
+H-STAR employs semantic reasoning universally, using it exclusively for direct lookup, common-sense, and complex lexical queries while using an additional SQL step for quantitative, mathematical, and logical tasks.
+This strategy optimizes performance on diverse question types, supporting text comprehension with computation when required.
+We demonstrate the efficacy of H-STAR across three tabular benchmarks involving table QA and table fact verification.
 Our experiments show that using the combined approach of H-STAR, aligned with multi-view extraction and adaptive reasoning outperforms all prior state-of-the-art approaches.
-Our results demonstrate that H-STAR is robust and generalizable across LLMs.
-Through ablations, error analysis, and performance evaluations, we illustrate the effectiveness of H-STAR.
+Our analysis demonstrates that H-STAR is robust and generalizable across LLMs, showcasing its superior reasoning capabilities across various tasks.
 Our main contributions are:
@@
 (b) Through extensive experiments, we demonstrate that H-STAR’s hybrid approach outperforms individually implemented symbolic and semantic methods, particularly excelling on longer tables.
-(c) Our experiments reveal H-STAR’s efficiency in table extraction using its precise ‘multi-view’ mechanism.
-Our analysis also highlights H-STAR’s superior reasoning capabilities, emphasizing the necessity of every step.
+(c) Our experiments showcase H-STAR’s efficient table extraction and enhanced reasoning capabilities across different models and datasets, emphasizing the importance of every stage in achieving improved performance.
 2 H-STAR Approach
-[…]
+H-STAR decomposes the table reasoning task into extraction and reasoning stages, combining LLM’s textual comprehension with symbolic reasoning.
+It converts the original table (T) into a query-specific table (TCR) (refer Algorithm 1).
+Figure 3 illustrates our H-STAR approach.
+Unlike, DATER Ye et al.
+(2023) and TabQSLify Nahid and Rafiei (2024), which use textual reasoning and text-to-SQL respectively, H-STAR integrates both reasoning types in a complementary manner for table extraction.
+Our reasoning step dynamically uses symbolic techniques alongside semantic methods for quantitative questions, overcoming the limits of pure textual reasoning.
+H-STAR operates in two main stages: 1) Table Extraction and 2) Adaptive Reasoning.
 2.1 Table Extraction
-Two-dimensional tables consist of columns (attributes) and rows (values).
+Two-dimensional tables consist of columns and rows.
 Table extraction involves a two-step reasoning chain: (1) column extraction and (2) row extraction (see Figure 3).
-As illustrated in Algorithm 1, for a given table T, table extraction returns a table Tcr (Steps 1-8) based on the input query Q.
+As illustrated in Algorithm 1, for a given table T, table extraction returns a table TCR (lines 1-8) based on the input query Q.
 Figure 3: An overview of our proposed method, H-STAR, consisting of a combination of code generation and text-based verification.
@@
 Column Extraction.
-As shown in Figure 3, H-STAR uses a ‘multi-view’ technique for column extraction, working with both the original table T (Step 1) and its transposed form TT (Step 2).
-This ensures that the LLM semantically understands and extracts correct attributes regardless of the table structure.
-Given an input table T and the question Q (‘How long did it take the New York Americans to win the National Cup after 1936?’), the LLM employs colsql to generate a SQL query to extract columns ‘year’ (C1).
-This is followed by the text-based column verification step coltext on TT returning a new set of columns C2 (‘year’, ‘national cup’).
-Table T is filtered for columns ‘year’, ‘national cup’ (C) derived from Steps 1 and 2 to obtain a column-filtered table $\textsc{T}_{\textsc{c}}$ (Step 4).
+As shown in Algorithm 1, H-STAR uses a ‘multi-view’ technique, first using the original table T (line 1) followed by its transposed form TT (line 2) for column extraction.
+Changing the ‘view’ acts as a verifier accounting for any information missed due to the table structure, thus improving the overall performance (see Appendix A.4).
+In Figure 3, given an input table T and the question Q (‘How long did it take the New York Americans to win the National Cup after 1936?’), the LLM employs colsql to generate an SQL query to extract columns ‘year’ (C1).
+This is followed by the text-based step coltext which processes table TT to return columns C2 (‘year’, ‘national cup’).
+This step captures the ‘national cup’ information relevant to answer the query, which is missed by colsql.
+Table T is then filtered for columns ‘year’, ‘national cup’ (C’) derived from C1 and C2 to obtain table Tc (lines 3,4).
 Algorithm 1 H-STAR
 0:  (T, Q) $\rightarrow$ Table - Question pair
-0:  $\hat{A}$rightarrow Predicted AnswerTable ExtractionStep 1: Column Extraction
-1:   C1$\leftarrow$ colsql(T, Q) $\triangleright$ SQL extraction
-2:   C2 $\leftarrow$ coltext(TT, Q, C1) $\triangleright$ Text $\triangleright$ $\textsc{T}^{\textsc{t}}$ refers to the transposed table
+0:  $\hat{A}$rightarrow Predicted AnswerStage A: Table ExtractionColumn Extraction
+1:   C1$\leftarrow$ colsql(T, Q) $\triangleright$ SQL approach
+2:   C2 $\leftarrow$ coltext(TT, Q) $\triangleright$ Text approach $\triangleright$ $\textsc{T}^{\textsc{t}}$ refers to the transposed table
 […]
@@
 7:   ${}^{\prime}\leftarrow$ R1 $\cup$ R2
-8:   TCR $\leftarrow$ TC.filter(R) $\triangleright$ Filter table $\textsc{T}_{\textsc{c}}$ with rows $\textsc{R}^{\prime}$ to get $\textsc{T}_{\textsc{cr}}$daptive Reasoning
+8:   TCR $\leftarrow$ TC.filter(R) $\triangleright$ Filter table $\textsc{T}_{\textsc{c}}$ with rows $\textsc{R}^{\prime}$ to get $\textsc{T}_{\textsc{cr}}$tage B: Adaptive Reasoning
 9:  if math(Q) ==absent$=$ = True then
@@
 Row Extraction.
-After obtaining the filtered table $\textsc{T}_{\textsc{c}}$ with relevant columns, H-STAR proceeds to the row extraction phase culminating in the query-specific table Tcr (Steps 5-8).
-The filtered table $\textsc{T}_{\textsc{c}}$ has fewer tokens than T, enabling more efficient row extraction.
-The SQL query generated by rowsql on the column-filtered table $\textsc{T}_{\textsc{c}}$ extracts the relevant rows R1 (‘row 18’).
+After obtaining the filtered table TC with relevant columns (Algorithm 1, line 4), H-STAR proceeds to the row extraction phase culminating in the query-specific table TCR (lines 5-8).
+The filtered table TC has fewer tokens than T, fitting within the token limit and enabling more efficient row extraction.
+As shown in Figure 3, the SQL query generated by rowsql on the column-filtered table TC extracts the relevant rows R1 (‘row 18’).
 While it returns the appropriate row, it is not sufficient to answer the question.
-To overcome these limitations, we use text-based row verification rowtext to obtain the missing rows, ultimately retrieving ‘row 1’ and ‘row 18’ ($\textsc{R}_{2}$).
-Combining $\textsc{R}_{1}$ and $\textsc{R}_{2}$ gives the relevant rows $\textsc{R}^{\prime}$ (‘row 1’, ‘row 18’), resulting in the final table $\textsc{T}_{\textsc{cr}}$.
+To overcome these limitations, we use text-based row verification rowtext to obtain the missing rows, ultimately retrieving ‘row 1’ and ‘row 18’ (R2).
+Combining R1 and R2 gives the relevant rows R’ (‘row 1’, ‘row 18’), resulting in the final table TCR.
 2.2 Adaptive Reasoning
-To leverage the strengths of textual reasoning while overcoming its quantitative limitations, we introduce an adaptive reasoning framework.
-Assessing the nature of the question, the LLM incorporates symbolic reasoning only when required.
-For instance, in Figure 3, when faced with a question like "How long did it take the New York Americans to win the National Cup after 1936?", LLM requires a count as an answer, and our pipeline prioritizes symbolic reasoning (fsql) through SQL-generated code.
-This code is executed on a SQL engine, enabling precise answers from the table $\textsc{T}_{\textsc{cr}}$.
-The output from this SQL-based evaluation Ev serves as additional evidence, enhancing the textual reasoning process (ftext).
-By explicitly integrating this quantitative approach, our algorithm effectively addresses LLMs’ limitations in handling mathematical operations, thereby improving the accuracy and robustness when tackling questions that demand numerical reasoning.
+We propose an adaptive reasoning framework that harnesses the strengths of textual reasoning while mitigating its quantitative limitations through symbolic reasoning.
+First, we prompt LLMs to leverage their language understanding capabilities to evaluate the query requirements.
+We apply symbolic reasoning for queries necessitating quantitative analysis and use the generated output (Ev) along with the table (TCR) as the input to the final textual reasoning step.
+For instance, in Figure 3, when faced with a question like "How long did it take the New York Americans to win the National Cup after 1936?", LLM needs to calculate the answer, and our pipeline prioritizes symbolic reasoning (fsql) through SQL-generated code.
+This code is executed on a SQL engine, enabling precise answers from the table TCR.
+The output from this SQL-based evaluation Ev serves as additional evidence, supporting the final reasoning step (ftext).
+By explicitly integrating this quantitative approach, our algorithm effectively addresses LLMs’ limitations in handling quantitative reasoning, thereby improving the accuracy and robustness when tackling questions that demand numerical reasoning.
 3 Experiments
@@
 LLM Models.
-In our research, we use state-of-the-art large language models (LLM) such as PaLM-2 Anil et al.
-(2023) and GPT-3.5-Turbo Brown et al.
-(2020); OpenAI (2023) for table reasoning tasks.
+In our research, we use state-of-the-art large language models (LLM) such as Gemini-1.5-Flash Reid et al.
+(2024), PaLM-2 Anil et al.
+(2023), GPT-3.5-Turbo, GPT-4o-mini OpenAI (2023), and the open-source Llama-3-70B Dubey et al.
+(2024) for table reasoning tasks.
 Our model inputs include in-context examples, the table, and the question for each step of the pipeline.
-Previous studies, such as Chain-of-Table and TabSQLify, influenced our model choices to ensure a fair comparison.
 All baselines with PALM-2 are from the propriety PALM-2 model, which is different from the public PALM-2 API (used in H-STAR and TabSQLify).
@@
 TabSQLify uses SQL to trim the table and then reasons based on the reduced table.
-3.1 Results and Analysis
-Table 1 compares the performance of different methods on the table reasoning tasks, TabFact, and WikiTQ datasets, across several LLMs including GPT-3.5-Turbo and PaLM-2.
-This comparison involves evaluating against generic reasoning, table manipulation techniques, and our method, H-STAR.
-Appendix A provides results on the FeTaQA dataset in addition to comparisons with other baseline models on the WikiTQ and TabFact dataset.
+3.1 Main Results
+Table 1 compares the performance of different methods on the TabFact, and WikiTQ datasets, across GPT-3.5-Turbo and PaLM-2.
+This comparison involves evaluating against generic reasoning, table manipulation techniques, and H-STAR.
+Appendix A includes analyses on the FeTaQA dataset.
 GPT-3.5-Turbo PaLM-2
@@
 TabSQLify 79.50 64.70 79.78 55.78
-H-STAR 83.74 68.85 85.92 68.51
+H-STAR 85.03 69.56 86.51 68.62
 Table 1: Comparison of various methods across datasets.
 The results are reported from the Chain-of-Table paper for fair comparison.
-* Chain-of-Table uses a proprietary PaLM-2 model that is better than their publicly available one.
+* Chain-of-Table uses a proprietary PaLM-2 model that is better than the publicly available version.
 Analysis.
-(a) On the WikiTQ dataset, H-STAR demonstrates superior performance with both GPT-3.5-Turbo and PaLM-2 models, surpassing all baseline methods.
-Specifically, H-STAR achieves an accuracy of 68.85% on WikiTQ with GPT-3.5-Turbo, marking a significant improvement of 17.01% over the vanilla GPT-3.5-Turbo model and a 4% increase over the previous best method, TabSQLify.
-(b) On the TabFact dataset, H-STAR outperforms all methods with GPT-3.5-Turbo, attaining an accuracy of 83.74%, which is a 13.29% improvement over the vanilla model and a 4% margin over the previous best, Chain-of-Table.
-However, with PaLM-2, H-STAR achieves comparable performance on TabFact while outperforming Chain-of-Table by 1.2% on WikiTQ.
+(a) On WikiTQ dataset, H-STAR surpasses all baselines with both GPT-3.5-Turbo and PaLM-2 models.
+H-STAR achieves an accuracy of 69.56% on WikiTQ with GPT-3.5-Turbo, marking an improvement of 17.72% over the vanilla GPT-3.5-Turbo model.
+(b) On the TabFact dataset, H-STAR outperforms all methods with GPT-3.5-Turbo, attaining an accuracy of 85.03%, which is a 14.58% improvement over the vanilla model.
+With PaLM-2, H-STAR achieves comparable performance with Chain-of-Table on TabFact while improving over it by 1.3% on WikiTQ.
+Comparison Across Methods
+Table 2 compares H-STAR with ReAcTable Zhang et al.
+(2023) and SYNTQA Zhang et al.
+(2024b), which use both textual and symbolic approaches for table reasoning.
+SYNTQA (GPT) uses the GPT model to decide between semantic and symbolic approaches, applying the chosen method for reasoning based on its selection.
+Additionally, it is compared with ALTER Zhang et al.
+(2024a), which augments both queries and table data to facilitate reasoning.
+TabFact WikiTQ
+ReAcTable 73.1 52.5
+ALTER 84.3 67.4
+SYNTQA (GPT) - 65.2
+H-STAR 85.0 69.6
+Table 2: Comparison of various methods across datasets on GPT-3.5-Turbo.
+Analysis.
+H-STAR outperforms methods that use symbolic and textual approaches, such as ReAcTable, SYNTQA(GPT), and ALTER, showing the effectiveness of its complementary reasoning in a two-stage process.
+Appendix A.1 presents a comparison with more methods.
+Performance Across LLMs
+Table 3 compares the performance of more advanced models like Gemini-1.5-Flash, GPT-4o-mini, and the open-source Llama-3-70B on TabFact and WikiTQ datasets.
+The results emphasize H-STAR’s generalizability across different models and consistent improvements across datasets.
+Even advanced models show significant benefits by using H-STAR.
+GPT-4o-mini Gemini-1.5 Llama-3
+TF WTQ TF WTQ TF WTQ
+Generic Reasoning
+End-to-End QA 73.22 59.43 81.12 58.47 78.41 57.89
+CoT 75.99 64.31 79.99 64.11 75.34 65.49
+Table Manipulation
+TabSQLify 78.30 68.74 79.50 63.92 60.70 66.85
+Chain-of-Table 85.09 68.53 86.95 70.05 85.86 70.76
+H-STAR 89.42 74.93 89.08 73.14 89.23 75.76
+Table 3: Comparison of various models across datasets. TF: TabFact; WTQ: WikiTQ
+Analysis.
+(a) H-STAR shows much higher accuracy for GPT-4o-mini, scoring 89.42% on TabFact and 74.93% on WikiTQ, compared to the End-to-End QA and Chain of Thought methods, which achieve 73.22% and 75.99% on TabFact and 59.43% and 64.31% on WikiTQ.
+This marks a 16.2% improvement on TabFact and 15.5% on WikiTQ.
+(b) For Gemini-1.5-Flash, accuracy rises from 81.12% and 58.47% respectively to 89.08% on TabFact and 73.14% on WikiTQ.
+(c) Similarly, for Llama-3-70B, H-STAR improves the performance across both the datasets by a margin of 11% on TabFact and 18% on WikiTQ, showing consistent gains across models and datasets.
+3.2 Efficiency Analysis
 Efficiency of Table Extraction.
@@
 Our table extraction method uses a two-step chain to select the relevant columns followed by the rows.
-Figure 4 showcases the effectiveness of the technique to extract only the relevant information from the table.
 Figure 4 compares the number of average table cells for the extracted table for H-STAR, H-STAR without row extraction, and H-STAR without column extraction with other baselines.
+Figure 4: Comparison of average table cells in the final table.
 Analysis.
@@
 This emphasizes how H-STAR effectively reduces irrelevant information, thus focusing LLMs on the right evidence for right reasoning.
-Figure 4: Comparison of average table cells in the final table.
-Performance on Longer Tables.
+Efficiency on Longer Tables.
 Table-CoT Chen (2023) highlights that longer table size, with tables exceeding 30 rows, is a significant cause of erroneous generations.
 This observation is supported by a decline in LLM performance as the number of tokens in the table increases.
-We classify LLM performance based on total tokens into three groups - small (< 2000 tokens), medium (2000 to 4000 tokens), and large (> 4000 tokens).
-We then compare H-STAR with other table manipulation methods, including BINDER, DATER, Chain-of-Table, and TabSQLify.
+In Table 4, we classify LLM performance based on total tokens into three groups - small (< 2000 tokens), medium (2000 to 4000 tokens), and large (> 4000 tokens) and compare H-STAR with table manipulation like methods BINDER, DATER, Chain-of-Table, and TabSQLify.
 Method Small Medium Large
@@
 TabSQLify 68.15 57.91 52.34
-H-STAR 71.19 64.07 62.50
+H-STAR 71.64 65.20 64.84
 Table 4: Performance of various methods on different table sizes on WikiTQ.
@@
 While other table manipulation methods experience a drop in performance with larger tables, H-STAR maintains consistent performance across all table sizes.
-Figure 5: Comparison of SQL-based and text-based column extraction methods for answering table queries: (a) SQL-based selection accurately identifies the columns ("1940/41", "description losses"), while text-based selection mistakenly includes "murdered" as a table column.
-(b) The text-based method selects both required columns ("year" and "national cup"), while the SQL-based approach overlooks "national cup" as a requirement.
-Figure 6: Comparison of SQL-based and text-based row extraction methods for answering table queries.
-(a) SQL-based selection accurately selects the necessary rows ("Murdered"), while text-based selection incorrectly includes additional rows; (b) Text-based method correctly selects relevant rows ("1953/54", "1936/37"), while SQL-based selection misinterprets the query, including only the row for the year "1953/54".
-3.2 Efficiency Analysis
-In Table 3, we study H-STAR by considering the total number of samples generated by the LLMs.
-BINDER and DATER utilize self-consistency techniques to refine their results, while the Chain-of-Table approach adopts a more iterative sample creation process.
-Specifically, BINDER generates 50 Neural-SQL samples, implementing self-consistency, while DATER employs self-consistency at each step, resulting in 100 samples.
-In contrast, the Chain-of-Table method opts for a more resource-efficient strategy, producing 25 samples across three steps: ‘Dynamic Plan’, ‘Generate Args’, and ‘Query’.
-TabSQLify generates the fewest samples, with one generation each for the table decomposition and query steps.
-Using single outputs can result in a decrease in accuracy when the model fails to return an answer.
-Thus, H-STAR incorporates self-consistency measures by generating two outputs at each stage to ensure output validity.
-Each step, except Query, involves two generations for both SQL and Text, i.e., a total of four steps for column and row retrieval.
-These additional steps serve as a precaution against abnormal LLM generations.
-Overall, H-STAR is efficient and only uses 6-10 sample generations.
-Method # samples / step Total # samples
-BINDER Neural SQL: 50 50
-DATER Decompose Table: 40 100
-Generate Cloze: 20
-Generate SQL: 20
-Query: 20
-Chain-of-Table Dynamic Plan $\leq$ 5 $\leq$ 25
-Generate Args $\leq$ 19
-Query: 1
-TabSQLify Table Decompose: 1 2
-Query: 1
-H-STAR Column Retrieval: 2-4 6-10
-Row Retrieval: 2-4
-Query: 2
-Table 3: Number of generated samples for different methods.
+3.3 Error Analysis
+H-STAR
+The disjoint, step-wise nature of H-STAR enables identifying and analyzing failures.
+Figure 5 shows 100 randomly selected instances from the Tabfact and WikiTQ datasets where H-STAR returns incorrect answers.
+In this study, ‘Missing Columns’ and ‘Missing Rows’ refer to missing necessary columns and rows, respectively.
+‘Incorrect Reasoning’ occurs when H-STAR extracts the correct table, but LLM fails to produce the correct answer.
+‘Incorrect Annotations’ include semantically identical answers in different formats, ambiguous questions, and incorrect gold answers.
+Figure 5 shows that for 100 TabFact failures, 2% are missing columns, 9% missing rows, 79% incorrect LLM reasoning, and 10% incorrect annotations.
+For WikiTQ, 6% are missing columns, 17% missing rows, 54% incorrect reasoning, and 23% incorrect annotations.
+Figure 5: Error distribution on 100 error samples across datasets for H-STAR (GPT-3.5-Turbo).
+Analysis.
+Fewer errors in column and row extraction indicate that H-STAR effectively retrieves the necessary table data.
+This improvement allows us to "shift left" by addressing issues earlier in the pipeline, enhancing overall performance.
+The higher percentage of errors in reasoning does not reflect poorly on the LLM; instead, it underscores the effectiveness of our table extraction process.
+H-STAR vs Others
+Figure 6 compares errors in H-STAR, with TabSQLify, and BINDER on 100 WikiTQ samples where TabSQLify fails.
+TabSQLify extracts wrong tables for 62 out of 100 (6 missing columns, 56 missing rows).
+Out of the 38 samples that remain, it incorrectly reasons on 29 of the samples (76%).
+BINDER does not handle table extraction and instead relies on generating multiple neural SQL queries for final reasoning.
+However, it fails to reason correctly in 70 out of 100 instances.
+Figure 6: Analysis of error types in 100 samples from WikiTQ where TabSQLify fails. Note: As we move from left to right, the total samples decrease for each stage in the pipeline.
+In contrast, H-STAR proves more effective in both table extraction and reasoning, generating correct answers on 43 samples.
+Additionally, it demonstrates fewer errors in table extraction (4 missed columns; 16 missed rows) and reasoning (27 of the remaining 80 samples, i.e.
+34%).
 4 Ablation Study
-We conduct an ablation study to measure the contribution of each step in the pipeline and assess the impact of integrating symbolic-textual and adaptive reasoning in H-STAR.
-4.1 Importance of Each Stage
-To evaluate the significance of the two primary stages in H-STAR: a) Table extraction and b) Adaptive reasoning, we conducted an ablation study.
-This study involved systematically removing one stage at a time.
+To evaluate the significance of the two primary stages in H-STAR: a) Table extraction and b) Adaptive reasoning, we conducted an ablation study, removing one stage at a time.
 For the first condition, the final adaptive reasoning stage was substituted with a basic chain-of-thought reasoning process, while maintaining the table extraction stage.
@@
 It highlights the importance of each step in the pipeline.
-Removing any step can result in a performance loss except for a marginal increase in the case of row extraction for TabFact.
+Removing any step can result in a performance loss especially adaptive reasoning which causes a 7% drop for both TabFact and WikiTQ datasets.
+For a detailed error case study and further analysis, refer to the Appendix E.
 Method TabFact WikiTQ
-H-STAR 85.92 68.51
+H-STAR 86.51 68.62
 w/o row extraction 86.17 66.30
@@
 Table 5: Performance on TabFact and WikiTQ datasets.
-4.2 Impact of a Hybrid SQL-Text Approach
-We perform a quantitative evaluation of the impact of the hybrid approach in H-STAR.
-Figures 5, 6 depict how using only one of the approaches results in insufficient outputs.
-Furthermore, it is evident from Table 5 that not employing a hybrid approach (H-STAR) results in a significant performance loss.
+4.1 Impact of the Hybrid Approach
+We perform a quantitative study of the impact of the hybrid approach in H-STAR.
+Figures 16, 17 (Appendix E) depict how using only one of the approaches results in insufficient outputs for column and row extraction tasks respectively.
+Furthermore, Table 6 shows that not employing a hybrid approach results in a significant performance loss.
 Method TabFact WikiTQ
@@
 H-STAR ${}_{\text{sql}}$ 58.50 46.09
-H-STAR 85.93 68.51
+H-STAR 86.51 68.62
 Table 6: Performance using only text-based/SQL-based reasoning; H-STAR text: H-STAR with textual reasoning only; H-STAR sql: H-STAR with symbolic reasoning only.
-The low accuracy of SQL approaches can be attributed to the incorrect final answers.
-The text-based reasoning approach achieves reasonably high scores on TabFact and WikiTQ, albeit lower than the hybrid approach.
-By combining textual and structured reasoning capabilities, the hybrid approach i.e., H-STAR can effectively handle diverse data formats and task requirements, leading to superior performance across various benchmarks.
-5 Where do LLMs fail?
-This section examines errors in H-STAR, contrasts the failure points with existing approaches, and summarizes our key findings.
-5.1 Error Analysis on H-STAR
-The disjoint, step-wise nature of H-STAR enables identifying and analyzing failures.
-Table 6 shows 100 randomly selected incorrect answer instances from the Tabfact and WikiTQ datasets.
-In this study, ‘Missing Columns’ and ‘Missing Rows’ refer to missing necessary columns and rows, respectively.
-‘Incorrect Reasoning’ occurs when H-STAR extracts the correct table, but LLM fails to produce the correct answer.
-‘Incorrect Annotations’ include semantically identical answers in different formats, ambiguous questions, and incorrect gold answers.
-Table 6 shows that for 100 TabFact failures, 2% are missing columns, 9% missing rows, 79% incorrect LLM reasoning, and 10% incorrect annotations.
-For WikiTQ, 6% are missing columns, 17% missing rows, 54% incorrect reasoning, and 23% incorrect annotations.
-Tabfact WikiTQ
-Missing Columns 2% 6%
-Missing Rows 9% 17%
-Incorrect Reasoning 79% 54%
-Incorrect Annotation 10% 23%
-Table 6: Error analysis on 100 error samples across datasets showcasing the distribution of errors for H-STAR.
 Analysis.
-Fewer errors in column and row extraction indicate that H-STAR efficiently extracts the necessary table.
-By reducing errors in the table extraction process, we effectively "shift left" addressing problems earlier in the pipeline and enhancing overall performance.
-5.2 More Error Analysis: H-STAR vs Others
-Table 7 compares errors in H-STAR, with TabSQLify, and BINDER.
-We use 100 WikiTQ samples where TabSQLify fails.
-Many errors with TabSQLify (6/100 missing columns, 56/94 missing rows) result from incorrect row extraction.
-Out of 100, TabSQLify correctly extracts the table for 38 samples, but incorrectly reasons on 29 (76%).
-BINDER does not extract tables but generates multiple neural SQL queries for table reasoning tasks.
-TabSQLify BINDER H-STAR
-Missing Columns 6/100 N/A 4/100
-Missing Rows 56/94 N/A 16/96
-Incorrect Reasoning 29/38 70/100 27/80
-Incorrect Annotation 9/9 9/30 10/53
-Correct Answers 0/0 21/21 43/43
-Table 7: Analysis of error types in 100 samples from WikiTQ where TabSQLify fails.
-Note: The denominator refers to the remaining samples; we start with 100 samples and subtract the number of error samples to determine the remaining samples.
-However, its output shows that it fails to reason correctly in 70 out of 100 cases.
-In contrast, H-STAR proves more effective in both table extraction and reasoning, achieving correct generations in 43 samples.
-Additionally, it demonstrates fewer errors in table extraction (4/100 columns; 16/96 rows) and reasoning (27/80, i.e., 34% of samples).
-5.3 What did we learn?
+SQL methods though efficient with numerical reasoning, are highly sensitive to data variations like irregular formatting and mixed data types, leading to a performance drop.
+In contrast, textual reasoning is more resilient, offering fuzzy matching and better interpretation of data.
+By combining textual and symbolic reasoning, H-STAR uses their complementary strengths, achieving superior performance.
+4.2 Symbolic vs Semantic Approaches
+As shown in Table 7 removing SQL extraction drops the accuracy to 85.22% on TabFact and 64.39% on WikiTQ, removing text extraction causes a drop to 83.74% and 60.31%, respectively.
+SQL reasoning removal results in 84.48% and 64.76%, but omitting text reasoning causes the largest drop, to 58.70% and 54.35%.
+Method TabFact WikiTQ
+H-STAR 86.51 68.62
+w/o SQL extraction 85.22 64.39
+w/o text extraction 83.74 60.31
+w/o SQL reasoning 84.48 64.76
+w/o text reasoning 58.70 54.35
+Table 7: Performance of H-STAR after systematically removing symbolic part and semantic parts from: 1) table extraction (both row and column extraction); 2) adaptive reasoning.
+Analysis.
+Table 7 shows that text-based approaches excel in both table extraction and reasoning.
+The evaluation datasets often contain noisy data that SQL-based approaches struggle with due to their reliance on structured schemas.
+Since text-based reasoning methods are more effective at handling such irregularities, they achieve reasonably high scores, albeit lower than our hybrid approach.
+5 Key Findings
 Firstly, our evaluation demonstrates that our hybrid approach achieves substantial improvements, surpassing the performance of previous state-of-the-art methods across various table reasoning tasks.
-Secondly, the quantitative analysis demonstrates that our ‘multi-view’ approach extracts tables with reduced irrelevant data, resulting in more concise tables.
-Furthermore, the qualitative analysis highlights fewer errors in our table extraction method compared to prior approaches, confirming a decrease in extraneous information.
+Secondly, the quantitative analysis demonstrates that our ‘multi-view’ approach extracts tables specific to the query.
+Furthermore, the qualitative analysis highlights fewer errors in our table extraction method compared to prior approaches, confirming a decrease in irrelevant information.
 Thirdly, our analysis indicates consistent performance even with longer tables, emphasizing our method’s effectiveness in accurately extracting relevant information by filtering out noise.
-The adaptive reasoning component boosts performance, especially for longer tables.
-Lastly, the ablation study shows that breaking down the task into sub-tasks significantly enhances the overall performance, with each sub-task playing a crucial role in achieving superior results.
-Moreover, it highlights the constraints of relying solely on either text or SQL approaches, which are effectively addressed by our H-STAR approach.
+Lastly, the ablation study shows that decomposing the task into sub-tasks significantly enhances the overall performance, with each sub-task playing a crucial role in achieving superior results.
+Moreover, it also highlights the constraints of relying solely on either text or SQL approaches, which are effectively addressed by our H-STAR approach.
 Together, these findings emphasize the substantial advantages of our hybrid and multi-view approach in addressing complex table reasoning tasks.
@@
 (2023); Nahid and Rafiei (2024) or by employing iterative methods Jiang et al.
+(2023); Zhang et al.
 (2023); Wang et al.
@@
 (2022) is an SQL-based approach that modifies SQL statements to include LLM API calls within SQL statements.
+ALTER Zhang et al.
+(2024a) augments both the queries along with the table data.
 DATER Ye et al.
@@
 Chain-of-Table Wang et al.
-(2023) uses a textual reasoning approach to update tables iteratively.
-The final table post-update is used for final reasoning.
-Unlike H-STAR, these techniques depend either on symbolic or semantic reasoning but not both.
+(2023) uses a textual reasoning approach to update tables iteratively before the final reasoning step.
+ReAcTable Zhang et al.
+(2023) extends the ReAct Yao et al.
+(2023) framework to table reasoning by employing step-by-step reasoning, iteratively generating sub-tables using external tools like SQL and Python.
+A detailed explanation has been provided in Appendix C.
+SYNTQA Zhang et al.
+(2024b) is an ensemble approach employing an answer selection mechanism selecting between text-to-SQL and text-based models.
+In contrast, H-STAR follows a fixed two-stage table extraction and reasoning pipeline, applying both symbolic and semantic approaches complementarily at each stage, which leads to improved performance.
 7 Conclusion
@@
 Lastly, we appreciate the reviewing team’s insightful comments.
-Appendix A Additional Results
-Tables 8 and 9 show results of other baseline models on WikiTQ and TabFact datasets respectively. Our method, H-STAR, outperformed all other models by substantial margins.
+Appendix A Supplementary Analysis
+A.1 Comparison With Diverse Methods
+Tables 8 and 9 show results of other baseline methods on WikiTQ and TabFact datasets respectively. Our method, H-STAR, outperformed all other models by substantial margins.
 Model Method Accuracy
-PLM TaPas 48.8
+Pre-trained/fine-tuned TaPas 48.8
 GraPPa 52.7
@@
 ITR 63.4
+SYNTQA(GPT)* 70.4
 Codex [^2] OpenAI Codex model is not available publically anymore. GPT-3 CoT 45.7
@@
 StructGPT 52.2
-GPT-3.5-Turbo H-STAR 68.8
+ALTER 67.4
+SYNTQA(GPT)* 65.2
+GPT-3.5-Turbo H-STAR 69.6
 Table 8: Comparison of H-STAR with additional methods on WikiTQ dataset.
+* SYNTQA (GPT) uses GPT-3.5-Turbo to choose between SQL/text reasoning and uses: 1) fine-tuned models for reasoning; 2) GPT-3.5-Turbo for reasoning.
 Model Method Accuracy
-PLM Table-BERT 68.1
+Pre-trained/fine-tuned Table-BERT 68.1
 LogicFactChecker 74.3
@@
 TableCoT 73.1
-GPT-3.5-Turbo H-STAR 83.7
+ALTER 84.3
+GPT-3.5-Turbo H-STAR 85.0
 Table 9: Comparison of H-STAR with additional methods on TabFact dataset.
+A.2 FeTaQA
+Represented in Table 10 are the results for the FeTaQA dataset using the ROUGE scores Lin (2004).
+We observe an incremental improvement in the scores when compared to other methods on the GPT-3.5-Turbo.
+The ROUGE scores focus on lexical similarities while ignoring the semantic similarity between predicted and gold outputs.
+These metrics are known to be insensitive to capturing improvements when using in-context learning since the model does not learn the expected style of the long-form text just from an instruction or a few examples.
 Prompting ROUGE-1 ROUGE-2 ROUGE-L
@@
 H-STAR 0.62 0.39 0.52
-Table 10: Performance on FeTaQA dataset. PaLM-2* refers to the Google proprietary PaLM-2 model
-Method Fluent Correct Adequate Faithful
-T5-large 94.6 54.8 50.4 50.4
-Human 95 92.4 95.6 95.6
-TableCoT 96 82 75 87
-Tabsqlify 97 88 84 93
-H-STAR 96.6 87.6 89.6 94
-Table 11: Human evaluation results on FeTaQA dataset.
-A.1 FeTaQA
-Represented in Table 10 are the results for the FeTaQA dataset using the ROUGE scores Lin (2004).
-We observe an incremental improvement in the scores when compared to other methods on the GPT-3.5-Turbo.
-The ROUGE scores focus on lexical similarities while ignoring the semantic similarity between predicted and gold outputs.
-These metrics are known to be insensitive to capturing improvements when using in-context learning since the model does not learn the expected style of the long-form text just from an instruction or a few examples.
+Table 10: Performance on FeTaQA dataset. PaLM-2* refers to the Google proprietary PaLM-2 model.
 Figure 7 illustrates an example where the generated output, despite being correct in answering the question, is penalized by the ROUGE metric.
@@
 While we marginally outperform other methods, there is still scope for improvement.
+Method Fluent Correct Adequate Faithful
+T5-large 94.6 54.8 50.4 50.4
+Human 95 92.4 95.6 95.6
+TableCoT 96 82 75 87
+TabSQLify 97 88 84 93
+H-STAR 96.6 87.6 89.6 94
+Table 11: Human evaluation results on FeTaQA dataset.
 Qualitative Analysis
@@
 This can be attributed to the use of the same LLM.
-Most of the failures for TabSQLify are derived from an insufficient table, which further highlights the limitations of using a single-view approach for table extraction.
-Figure 8 showcases such an example where TabSQLify returns ‘information not provided’ because of incorrect of an extracted table.
-It can be noted that H-STAR generates an answer comparable with the ground truth for the same.
-Figure 9 visually depicts the outputs from both methods.
-It illustrates the evaluation process undertaken to compare the quality of the outputs.
+Many failures for TabSQLify are derived from an insufficient table, which further highlights the limitations of using a single-view approach for table extraction.
 Method % of samples
@@
 Table 12: Comparison of H-STAR with TabSQLify on 100 samples.
+Figure 8 showcases such an example where TabSQLify returns ‘information not provided’ because of incorrect of an extracted table.
+It can be noted that H-STAR generates an answer comparable with the ground truth for the same.
+Figure 9 visually depicts the outputs from both methods.
+It illustrates the evaluation process undertaken to compare the quality of the outputs.
 Figure 8: An example of FeTaQA dataset where TabSQLify fails as a result of insufficient table evidence.
 Figure 9: Comparison of the outputs with TabSQLify.
+A.3 Analysis: Generated Samples
+In Table 13, we study H-STAR by considering the total number of samples generated by the LLMs.
+BINDER and DATER utilize self-consistency techniques to refine their results, while the Chain-of-Table approach adopts a more iterative sample creation process.
+Specifically, BINDER generates 50 Neural-SQL samples, implementing self-consistency, while DATER employs self-consistency at each step, resulting in 100 samples.
+In contrast, the Chain-of-Table method opts for a more resource-efficient strategy, producing 25 samples across three steps: ‘Dynamic Plan’, ‘Generate Args’, and ‘Query’.
+TabSQLify generates the fewest samples, with one generation each for the table decomposition and query steps.
+Using single outputs can result in a decrease in accuracy when the model fails to return an answer.
+Thus, H-STAR incorporates self-consistency measures by generating two outputs at each stage to ensure output validity.
+Each step, except Query, involves two generations for both SQL and Text, i.e., a total of four steps for column and row retrieval.
+These additional steps serve as a precaution against abnormal LLM generations.
+Overall, H-STAR is efficient and only uses 6-10 sample generations.
+Method # samples / step Total # samples
+BINDER Neural SQL: 50 50
+DATER Decompose Table: 40 100
+Generate Cloze: 20
+Generate SQL: 20
+Query: 20
+Chain-of-Table Dynamic Plan $\leq$ 5 $\leq$ 25
+Generate Args $\leq$ 19
+Query: 1
+TabSQLify Table Decompose: 1 2
+Query: 1
+H-STAR Column Extraction: 2-4 6-10
+Row Extraction: 2-4
+Query: 2
+Table 13: Number of generated samples for different methods.
+A.4 Benefit of using Transposed Table
+Table 14 compares H-STAR’s performance using transposed versus original tables for text-based column verification.
+On TabFact, H-STAR achieves 86.51% accuracy with transposed tables, versus 85.22% with original tables.
+On WikiTQ, accuracy is 68.62% with transposed tables, compared to 66.59% with original tables.
+Method TabFact WikiTQ
+H-STAR (with transposed table) 86.51 68.62
+H-STAR (with original table) 85.22 66.59
+Table 14: Comparison between H-STAR using a) transposed table and b) original table for text-based column extraction.
+Analysis.
+Transposing tables leads to better performance than using the original table for column verification.
+The transposed table approach is more effective for complex datasets like WikiTQ, which are harder to reason upon and contain longer tables.
 Appendix B Input Table Format
@@
 columns: [’name’, ’league’, ’total’]
-Appendix C Hyperparameters
-Table 13 provides the details of PaLM-2 hyper-parameters used for the WikiTQ and TabFact datasets whereas Table 14 showcases the hyperparameter setting for GPT-3.5-Turbo.
+Appendix C Comparison with ReAct
+While our method appears to be similar to ReAct in terms of combining textual reasoning and tool-based methods, there are significant differences between H-STAR and ReAct:
+1. Integration of Symbolic and Semantic Reasoning.
+H-STAR combines SQL-based and text-based reasoning, while ReAct merges textual reasoning with actions.
+ReAct uses tools mainly for actions, not reasoning.
+H-STAR applies SQL for quantitative tasks and text for lexical queries, whereas ReAct focuses more on textual reasoning.
+2. Fixed Multi-Stage Approach.
+ReAct prompts LLMs to alternate between reasoning and actions, allowing interaction with external sources.
+H-STAR, on the other hand, splits tabular reasoning into two stages: 1) Table Extraction, and 2) Adaptive Reasoning, blending SQL and text reasoning.
+This hybrid approach sets H-STAR apart from ReAct’s focus on textual reasoning and interaction with the environment.
+3. Table Extraction Approach.
+H-STAR uses a "multi-view" approach for table extraction, identifying relevant columns from both original and transposed tables before selecting rows, reducing irrelevant data, and improving focus.
+In contrast, ReAct does not employ any extraction method.
+4. Adaptive Reasoning Strategy.
+H-STAR uses SQL to support semantic reasoning for quantitative tasks while relying solely on semantic reasoning for lexical queries.
+In contrast, ReAct interleaves reasoning and actions without adapting to task type.
+H-STAR’s hybrid approach, combining SQL logic and text understanding, excels in handling diverse tasks, outperforming ReAct on table reasoning tasks.
+Appendix D Implementation Details
+We provide the implementation details including the hyperparameters and prompt details for each of the individual steps.
+H-STAR operates in two key stages: 1) Table Extraction and 2) Adaptive Reasoning.
+The Table Extraction phase includes column extraction (SQL-based: colsql and text-based: coltext) and row extraction (SQL-based: rowsql and text-based: rowtext).
+The Adaptive Reasoning phase consists of SQL-based (fsql) and text-based (ftext) reasoning components.
+D.1 Hyperparameters
+Table 15 provides the details of PaLM-2 hyper-parameters used for the WikiTQ and TabFact datasets.
+Table 16 showcases the hyperparameter setting for GPT-3.5-Turbo.
 The hyperparameters ‘samples’ indicate the number of outputs taken for each step of the pipeline whereas the ‘examples’ indicate the number of few-shot demonstrations for each step.
@@
 rowtext 0.7 1.0 512 2 2
-Reasonsql 0.1 1.0 512 1 3
-Reasontext 0.1 1.0 256 1 4
+fsql 0.1 1.0 512 1 3
+ftext 0.0 1.0 512 1 4
 TabFact
@@
 rowtext 0.7 1.0 512 2 3
-Reasonsql 0.1 1.0 512 1 3
-Reasontext 0.1 1.0 256 1 5
-Table 13: Hyperparameter settings for PaLM-2
+fsql 0.1 1.0 512 1 3
+ftext 0.0 1.0 512 1 5
+Table 15: Hyperparameter settings for PaLM-2.
 Function temperature top_p output_tokens samples examples
@@
 coltext 0.4 1.0 512 2 2
-Rowsql 0.4 1.0 512 2 3
+rowsql 0.3 1.0 512 2 3
 rowtext 0.4 1.0 512 2 3
-Reasonsql 0.1 1.0 512 1 4
-Reasontext 0.1 1.0 256 1 4
+fsql 0.1 1.0 512 1 4
+ftext 0.0 1.0 512 1 4
 TabFact
@@
 rowtext 0.5 1.0 512 2 3
-Reasonsql 0.1 1.0 512 1 4
-Reasontext 0.1 1.0 256 1 5
-Table 14: Hyperparameter settings for GPT-3.5-Turbo
+fsql 0.1 1.0 512 1 4
+ftext 0.0 1.0 512 1 5
+Table 16: Hyperparameter settings for GPT-3.5-Turbo.
+D.2 Input Prompts
+Figures 10, 11, 12, 13, 14 and 15 illustrate the input prompts containing sample in-context learning examples for the steps colsql, coltext, rowsql, rowtext, fsql, and ftext respectively.
+Figure 10: Prompt for SQL-based column extraction (colsql).
+Figure 11: Prompt for Column-based column verification (coltext).
+Figure 12: Prompt for SQL-based row selection (rowsql).
+Figure 13: Prompt for text-based row verification (rowtext).
+Figure 14: Prompt for SQL-based reasoning (fsql).
+Figure 15: Prompt for text-based reasoning (ftext).
+Appendix E Error Analysis: Case Study
+Figures 16 and 17 showcase how semantic and symbolic approaches can individually fail for the column and row extraction steps.
+Figures 19, 18, 20 and 21 illustrate error cases for each individual step within H-STAR.
+Figure 16: Comparison of SQL-based and text-based column extraction methods for answering table queries: (a) SQL-based selection accurately identifies the columns ("1940/41", "description losses"), while text-based selection mistakenly includes "murdered" as a table column.
+(b) The text-based method selects both required columns ("year" and "national cup"), while the SQL-based approach overlooks "national cup" as a requirement.
+Figure 17: Comparison of SQL-based and text-based row extraction methods for answering table queries.
+(a) SQL-based selection accurately selects the necessary rows ("Murdered"), while text-based selection incorrectly includes additional rows; (b) Text-based method correctly selects relevant rows ("1953/54", "1936/37"), while SQL-based selection misinterprets the query, including only the row for the year "1953/54".
+Figure 18: Case Study: ‘Missing Columns’.
+An illustration of the error ‘Missing Columns’.
+In the above example, the extracted table does not contain the column ‘Tournament’ that contains the ‘Year End Ranking’.
+The omission of the ‘Tournament’ column affects the downstream steps in the pipeline leading to an insufficient row extraction and reasoning.
+Figure 19: Case Study: ‘Missing Rows’.
+An illustration of the error ‘Missing Rows’.
+From the above example, it can be seen that while the extracted columns are correct, the row extraction is incorrect.
+Thus, as the extracted table lacks the necessary information, the final answer is incorrect.
+Figure 20: Case Study: ‘Incorrect Reasoning’.
+An illustration of the error ‘Incorrect Reasoning’.
+Despite the correct question-relevant table being extracted, the model misinterprets the question.
+It adds up the scores by the opposing team instead thus returning an incorrect answer.
+Figure 21: Case Study: ‘Incorrect Annotation’.
+An illustration of the error ‘Incorrect Annotation’.
+In the example, the query-specific table is correctly extracted and the final reasoning leads to the correct answer.
+However, the prediction is penalized for not being an ‘exact match’.

### v3

@@
 Our extensive experiments demonstrate that H-STAR significantly outperforms state-of-the-art methods across three tabular question-answering (QA) and fact-verification datasets, underscoring its effectiveness and efficiency.
-[^1] The codes and data are available at: https://github.com/nikhilsab/H-STAR
+[^1] The code and data are available at: https://github.com/nikhilsab/H-STAR
 1 Introduction
-Tabular data is one of the most widely used formats for storing structured information in real-world applications.
-Table-based reasoning presents an inherently challenging problem, requiring logical, mathematical, and textual reasoning over unstructured queries and structured tables Ye et al.
+Tabular data are among the most widely used formats for storing structured information in real-world applications.
+Tabular reasoning presents an inherently challenging problem, requiring logical, mathematical, and textual reasoning over unstructured queries and structured tables Ye et al.
 (2023).
 Thus, understanding and inferring tabular data has become a significant area of research in natural language processing.
-Tabular reasoning tasks (illustrated in Figure 1), such as table-based question answering Pasupat and Liang (2015); Nan et al.
+Tabular reasoning tasks (Figure 1), such as table-based question answering Pasupat and Liang (2015); Nan et al.
 (2022) and table-based fact verification Chen et al.
@@
 Figure 1: An illustration of different tabular reasoning tasks (a) Fact-verification, (b) Short-form QA, and (c) Long-form QA.
-For each task, a question Q is paired with its answer A, which varies by task.
+For each task, question Q is paired with its answer A, which varies by task.
 Evidence shows the relevant columns and rows needed to answer the question.
@@
 In the domain of table reasoning, symbolic approaches such as Program of Thought Chen et al.
-(2023) and textual methods like Chain of Thought (CoT) Wei et al.
+(2023) and textual methods such as Chain of Thought (CoT) Wei et al.
 (2022); Brown et al.
@@
 Textual reasoning excels in natural language understanding but often misinterprets table structures and struggles with quantitative reasoning.
-Conversely, SQL-based approaches are strong in quantitative problem-solving but perform poorly on noisy or unstructured inputs Liu et al.
+In contrast, SQL-based approaches are strong in quantitative problem solving, but perform poorly on noisy or unstructured inputs Liu et al.
 (2023).
@@
 (2023); Nahid and Rafiei (2024).
-Focusing on relevant table cells helps LLMs by providing the right context, thereby reducing hallucinations in reasoning.
+Focusing on relevant table cells helps LLMs by providing the right context, thus reducing hallucinations in reasoning.
 We employ a combination of SQL-based and text-based techniques in H-STAR’s table extraction phase.
 In the adaptive reasoning phase, we use the language comprehension capabilities of LLMs to guide them with few-shot examples to decide when to support semantic methods with symbolic (SQL) approaches.
-H-STAR employs semantic reasoning universally, using it exclusively for direct lookup, common-sense, and complex lexical queries while using an additional SQL step for quantitative, mathematical, and logical tasks.
+H-STAR employs semantic reasoning universally, using it exclusively for direct lookup, common sense, and complex lexical queries while using an additional SQL step for quantitative, mathematical, and logical tasks.
 This strategy optimizes performance on diverse question types, supporting text comprehension with computation when required.
@@
 Our analysis demonstrates that H-STAR is robust and generalizable across LLMs, showcasing its superior reasoning capabilities across various tasks.
-Our main contributions are:
+Our main contributions are as follows:
 (a) We introduce H-STAR, a two-step method unifying table extraction with adaptive reasoning to enhance performance on tabular reasoning.
 It is a hybrid approach that effectively integrates symbolic (SQL logic) and semantic (textual) methods.
-(b) Through extensive experiments, we demonstrate that H-STAR’s hybrid approach outperforms individually implemented symbolic and semantic methods, particularly excelling on longer tables.
-(c) Our experiments showcase H-STAR’s efficient table extraction and enhanced reasoning capabilities across different models and datasets, emphasizing the importance of every stage in achieving improved performance.
+(b) Through extensive experiments, we show that H-STAR’s hybrid approach surpasses state-of-the-art methods that rely solely on either symbolic or semantic techniques, especially when reasoning over larger tables.
+(c) Our ablation study and analysis highlight H-STAR’s effectiveness in table extraction and reasoning across various models and datasets, underscoring the significance of each stage in enhancing overall performance.
+The code and additional resources are available at: https://hstar-llm.github.io/.
 2 H-STAR Approach
@@
 Figure 3 illustrates our H-STAR approach.
-Unlike, DATER Ye et al.
+Unlike DATER Ye et al.
 (2023) and TabQSLify Nahid and Rafiei (2024), which use textual reasoning and text-to-SQL respectively, H-STAR integrates both reasoning types in a complementary manner for table extraction.
@@
 H-STAR operates in two main stages: 1) Table Extraction and 2) Adaptive Reasoning.
+Implementation details, including prompts, input table formats, and hyperparameters, are provided in Appendix C and D.
 2.1 Table Extraction
@@
 As illustrated in Algorithm 1, for a given table T, table extraction returns a table TCR (lines 1-8) based on the input query Q.
-Figure 3: An overview of our proposed method, H-STAR, consisting of a combination of code generation and text-based verification.
+Figure 3: An overview of H-STAR, consisting of a combination of code generation and text-based verification.
 Given a complex table and its question, H-STAR answers using (a) Table Extraction: extracts the question-specific table from the original by first selecting the columns followed by rows.
@@
 This is followed by the text-based step coltext which processes table TT to return columns C2 (‘year’, ‘national cup’).
-This step captures the ‘national cup’ information relevant to answer the query, which is missed by colsql.
+This step captures the ‘national cup’ information relevant to answering the query, which is missed by colsql.
 Table T is then filtered for columns ‘year’, ‘national cup’ (C’) derived from C1 and C2 to obtain table Tc (lines 3,4).
@@
 As shown in Figure 3, the SQL query generated by rowsql on the column-filtered table TC extracts the relevant rows R1 (‘row 18’).
-While it returns the appropriate row, it is not sufficient to answer the question.
+Although it returns the appropriate row, it is not sufficient to answer the question.
 To overcome these limitations, we use text-based row verification rowtext to obtain the missing rows, ultimately retrieving ‘row 1’ and ‘row 18’ (R2).
@@
 We apply symbolic reasoning for queries necessitating quantitative analysis and use the generated output (Ev) along with the table (TCR) as the input to the final textual reasoning step.
-For instance, in Figure 3, when faced with a question like "How long did it take the New York Americans to win the National Cup after 1936?", LLM needs to calculate the answer, and our pipeline prioritizes symbolic reasoning (fsql) through SQL-generated code.
+For instance, in Figure 3, when faced with a question like ‘How long did it take the New York Americans to win the National Cup after 1936?’, LLM needs to calculate the answer, and our pipeline prioritizes symbolic reasoning (fsql) through SQL-generated code.
 This code is executed on a SQL engine, enabling precise answers from the table TCR.
@@
 (2019): A fact verification benchmark utilizing Wikipedia tables.
-We evaluate the test-small set, containing 2,024 statements and 298 tables, with each statement labeled as Entailed ("True") or Refuted ("False").
+We evaluated the test-small set, containing 2,024 statements and 298 tables, with each statement labeled Entailed ("True") or Refuted ("False").
 (b) WikiTQ Pasupat and Liang (2015): The WikiTableQuestions (WikiTQ) dataset involves question-answering tasks over semi-structured Wikipedia tables.
@@
 We tailor our evaluation metrics based on the task and dataset characteristics.
-For fact-verification datasets like TabFact, we evaluate performance using binary classification accuracy.
+For fact verification data sets such as TabFact, we evaluated performance using binary classification accuracy.
 For short-form question-answering datasets such as WikiTQ, we assess accuracy by measuring the exact match between predicted outputs and the gold-standard answers.
-For more complex tasks in FeTaQA, which involve long-form question answering, we evaluate performance using ROUGE-1, ROUGE-2, and ROUGE-L metrics Lin (2004), comparing predicted outputs against the long-form answers.
+For more complex tasks in FeTaQA, which involve long-form question answering, we evaluate performance using ROUGE-1, ROUGE-2, and ROUGE-L metrics Lin (2004), comparing predicted outputs with long-form answers.
 LLM Models.
@@
 This comparison involves evaluating against generic reasoning, table manipulation techniques, and H-STAR.
-Appendix A includes analyses on the FeTaQA dataset.
+Appendix A.1 offers a comprehensive analysis of the FeTaQA dataset, including a comparison with baselines, human evaluation, and qualitative analysis.
 GPT-3.5-Turbo PaLM-2
@@
 With PaLM-2, H-STAR achieves comparable performance with Chain-of-Table on TabFact while improving over it by 1.3% on WikiTQ.
-Comparison Across Methods
+Comparison Across Methods.
 Table 2 compares H-STAR with ReAcTable Zhang et al.
@@
 Additionally, it is compared with ALTER Zhang et al.
-(2024a), which augments both queries and table data to facilitate reasoning.
+(2024a), which augments both queries and table data to facilitate reasoning and E5 Zhang et al.
+(2024c), leveraging code generation capabilities of LLMs for table extraction followed by reasoning.
 TabFact WikiTQ
 ReAcTable 73.1 52.5
+E5 75.6 50.9
 ALTER 84.3 67.4
@@
 Analysis.
-H-STAR outperforms methods that use symbolic and textual approaches, such as ReAcTable, SYNTQA(GPT), and ALTER, showing the effectiveness of its complementary reasoning in a two-stage process.
-Appendix A.1 presents a comparison with more methods.
-Performance Across LLMs
+H-STAR outperforms methods that use symbolic and textual approaches, such as ReAcTable, SYNTQA(GPT), E5, and ALTER, showing the effectiveness of its complementary reasoning in a two-stage process.
+Appendix A.2 presents a comprehensive comparison of methods, including pre-trained models, fine-tuned architectures, and state-of-the-art LLM-based approaches.
+Our evaluation demonstrates that H-STAR consistently achieves superior performance across all benchmark metrics compared to all the baselines.
+Performance Across LLMs.
 Table 3 compares the performance of more advanced models like Gemini-1.5-Flash, GPT-4o-mini, and the open-source Llama-3-70B on TabFact and WikiTQ datasets.
@@
 Analysis.
-In particular, H-STAR significantly reduces the average number of processed table cells - from 159 to 18 for the WikiTQ dataset, from 88 to 13 for the TabFact dataset, and from 86 to 10 for the FeTaQA dataset.
-In contrast, other table manipulation methods, like TabSQLify and DATER yield higher values.
-Our approach showcases superior effectiveness, notably reducing cell counts across all datasets.
-This emphasizes how H-STAR effectively reduces irrelevant information, thus focusing LLMs on the right evidence for right reasoning.
+H-STAR employs a structured, step-by-step process: it first extracts query-relevant columns before selecting relevant rows.
+This progressive filtering minimizes token usage at each stage, thereby reducing computational overhead.
+As a result, H-STAR significantly reduces the average number of processed table cells—dropping from 159 to 18 on the WikiTQ dataset, from 88 to 13 on TabFact, and from 86 to 10 on FeTaQA.
+In contrast, alternative table manipulation methods, such as TabSQLify and DATER, yield considerably higher values.
+By effectively filtering out irrelevant information, H-STAR enables LLMs to focus on the right evidence for right reasoning, demonstrating superior efficiency across all datasets.
 Efficiency on Longer Tables.
@@
 Analysis.
-The results shown in Table 4 emphasize the challenges LLMs face when reasoning over longer tables.
-Our findings align with previous research, revealing a significant decrease in LLM performance as table sizes increase.
-While other table manipulation methods experience a drop in performance with larger tables, H-STAR maintains consistent performance across all table sizes.
+The results presented in Table 4 highlight the challenges LLMs face when reasoning over longer tables.
+Our findings confirm previous research, showing a significant decline in LLM performance as table sizes increase.
+While other table manipulation methods suffer from this scaling issue, H-STAR maintains consistent performance across all table sizes.
+This success is attributed to H-STAR ’s efficient table extraction process, which reduces the table size by removing irrelevant data that would otherwise act as noise and hinder reasoning.
+Resource Efficiency.
+In Table 5, we analyze H-STAR by examining the total number of samples generated by LLMs.
+BINDER and DATER employ self-consistency techniques to refine their results, while Chain-of-Table follows an iterative sampling process.
+Specifically, BINDER generates 50 Neural-SQL samples using self-consistency, whereas DATER applies self-consistency at each step, producing 100 samples.
+In contrast, Chain-of-Table adopts a more resource-efficient approach, generating 25 samples across three steps: ‘Dynamic Plan’, ‘Generate Args’, and ‘Query’.
+TabSQLify generates the fewest samples, with a single generation for both table decomposition and query steps.
+However, relying on single outputs can reduce accuracy when the model fails to return a valid answer.
+To mitigate this, H-STAR integrates self-consistency by generating two outputs per stage, ensuring output validity.
+Each step, except ‘Query’, involves two generations for both SQL and text, totaling four steps for column and row retrieval.
+These additional steps act as safeguards against unreliable LLM outputs.
+Overall, H-STAR maintains efficiency, requiring only 6–10 sample generations.
+Method # samples / step Total # samples
+BINDER Neural SQL: 50 50
+DATER Decompose Table: 40 100
+Generate Cloze: 20
+Generate SQL: 20
+Query: 20
+Chain-of-Table Dynamic Plan $\leq$ 5 $\leq$ 25
+Generate Args $\leq$ 19
+Query: 1
+TabSQLify Table Decompose: 1 2
+Query: 1
+H-STAR Column Extraction: 2-4 6-10
+Row Extraction: 2-4
+Query: 2
+Table 5: Number of generated samples for different methods.
 3.3 Error Analysis
-H-STAR
+H-STAR.
 The disjoint, step-wise nature of H-STAR enables identifying and analyzing failures.
@@
 The higher percentage of errors in reasoning does not reflect poorly on the LLM; instead, it underscores the effectiveness of our table extraction process.
-H-STAR vs Others
+H-STAR vs Others.
 Figure 6 compares errors in H-STAR, with TabSQLify, and BINDER on 100 WikiTQ samples where TabSQLify fails.
@@
 Figure 6: Analysis of error types in 100 samples from WikiTQ where TabSQLify fails. Note: As we move from left to right, the total samples decrease for each stage in the pipeline.
+Analysis.
 In contrast, H-STAR proves more effective in both table extraction and reasoning, generating correct answers on 43 samples.
@@
 4 Ablation Study
-To evaluate the significance of the two primary stages in H-STAR: a) Table extraction and b) Adaptive reasoning, we conducted an ablation study, removing one stage at a time.
-For the first condition, the final adaptive reasoning stage was substituted with a basic chain-of-thought reasoning process, while maintaining the table extraction stage.
-Conversely, the second condition retained the adaptive reasoning stage but omitted the table extraction stage, specifically by removing the row extraction and column extraction steps individually and then collectively.
+To evaluate the significance of the two primary stages in H-STAR: (a) Table extraction and (b) Adaptive reasoning, we conducted an ablation study, removing one stage at a time.
+For the first condition, the final adaptive reasoning stage was substituted with a basic chain-of-thought reasoning process while maintaining the table extraction stage.
+Conversely, the second condition retained the adaptive reasoning stage but omitted the table extraction stage, specifically removing the row extraction and column extraction steps individually and then collectively.
 A detailed assessment of the contribution of each stage to the overall performance of the method is shown in Table 6.
@@
 We perform a quantitative study of the impact of the hybrid approach in H-STAR.
-Figures 16, 17 (Appendix E) depict how using only one of the approaches results in insufficient outputs for column and row extraction tasks respectively.
+Figures 16 and 17 (Appendix E) show how the use of only one of the approaches results in insufficient outputs for column and row extraction tasks, respectively.
 Furthermore, Table 7 shows that not employing a hybrid approach results in a significant performance loss.
@@
 Analysis.
-SQL methods though efficient with numerical reasoning, are highly sensitive to data variations like irregular formatting and mixed data types, leading to a performance drop.
+SQL methods though efficient with numerical reasoning, are highly sensitive to data variations such as irregular formatting and mixed data types, leading to a performance drop.
 In contrast, textual reasoning is more resilient, offering fuzzy matching and better interpretation of data.
@@
 4.2 Symbolic vs Semantic Approaches
-As shown in Table 7 removing SQL extraction drops the accuracy to 85.22% on TabFact and 64.39% on WikiTQ, removing text extraction causes a drop to 83.74% and 60.31%, respectively.
+As shown in Table 8 removing SQL extraction drops the accuracy to 85.22% on TabFact and 64.39% on WikiTQ, and removing text extraction causes a drop to 83.74% and 60.31%, respectively.
 SQL reasoning removal results in 84.48% and 64.76%, but omitting text reasoning causes the largest drop, to 58.70% and 54.35%.
@@
 w/o text reasoning 58.70 54.35
-Table 7: Performance of H-STAR after systematically removing symbolic part and semantic parts from: 1) table extraction (both row and column extraction); 2) adaptive reasoning.
+Table 8: Performance of H-STAR after systematically removing symbolic part and semantic parts from: (1) table extraction (both row and column extraction); (2) adaptive reasoning.
 Analysis.
@@
 (2022) work on pre-training language models jointly on large-scale tabular and text data to reason in an end-to-end manner.
-Advancements in LLMs allow them to learn from in-context examples, reducing inference costs.
+Advancements in LLMs, allow them to learn from in-context examples, reducing inference costs.
 Text-to-SQL Rajkumar et al.
@@
 (2023) use symbolic methods to solve table-based tasks via Python/SQL programs.
-Textual-based reasoning techniques like Table-CoT Chen (2023) and Tab-CoT Ziqi and Lu (2023) extend prompting methods such as zero and few-shot CoT for tabular reasoning.
-Decomposing problems into smaller, manageable tasks has proven effective in solving complex reasoning challenges Zhou et al.
+Textual-based reasoning techniques such as Table-CoT Chen (2023) and Tab-CoT Ziqi and Lu (2023) extend prompting methods such as zero- and few-shot CoT for tabular reasoning.
+The decomposition of problems into smaller and manageable tasks has proven effective in solving complex reasoning challenges Zhou et al.
 (2022); Khot et al.
@@
 DATER Ye et al.
-(2023) and TabSQLify Nahid and Rafiei (2024) are table decomposition methods that use semantic and symbolic reasoning respectively.
+(2023) and TabSQLify Nahid and Rafiei (2024) are table decomposition methods that use semantic and symbolic reasoning, respectively.
 Chain-of-Table Wang et al.
 (2023) uses a textual reasoning approach to update tables iteratively before the final reasoning step.
-ReAcTable Zhang et al.
-(2023) extends the ReAct Yao et al.
-(2023) framework to table reasoning by employing step-by-step reasoning, iteratively generating sub-tables using external tools like SQL and Python.
-A detailed explanation has been provided in Appendix C.
 SYNTQA Zhang et al.
-(2024b) is an ensemble approach employing an answer selection mechanism selecting between text-to-SQL and text-based models.
-In contrast, H-STAR follows a fixed two-stage table extraction and reasoning pipeline, applying both symbolic and semantic approaches complementarily at each stage, which leads to improved performance.
+(2024b) is an ensemble approach employing an answer selection mechanism that selects between text-to-SQL and text-based models.
+The works most closely related to our approach are ReAcTable Zhang et al.
+(2023) and E5 Zhang et al.
+(2024c).
+ReAcTable extends the ReAct framework Yao et al.
+(2023) to table reasoning by employing step-by-step reasoning, where it iteratively generates sub-tables using external tools like SQL and Python.
+Similarly, E5 introduces a multi-step process for hierarchical table question answering, interpreting table structures, generating and executing code, and performing reasoning over the results to derive answers.
+In contrast, H-STAR employs a fixed two-stage pipeline for table extraction and reasoning, utilizing both symbolic and semantic approaches complementarily at each stage, resulting in improved performance.
+A comprehensive comparison of the methodological differences and their implications for performance is provided in Appendix B.
 7 Conclusion
@@
 Limitations
-Our current work has focused primarily on a subset of table reasoning tasks using datasets sourced from Wikipedia.
+Our current work has primarily focused on a subset of table reasoning tasks using datasets sourced from Wikipedia.
 While this has laid a solid foundation, it limits exploration into diverse reasoning tasks such as table manipulation, text-to-table generation, and table augmentation, which could provide valuable insights and enhance our approach’s capabilities.
@@
 These structures present unique challenges that our current approach may not effectively address.
-Finally, our study has focused solely on the English language, potentially limiting applicability to languages with different linguistic complexities that we have not accounted for.
+Finally, our study has focused solely on the English language, potentially limiting its applicability to languages with different linguistic complexities that we have not accounted for.
 Ethics Statement
@@
 Acknowledgments
-We thank the authors of Tabsqlify and BINDER for sharing their code, and Varun Yerram for his help in setting up the implementation process.
-This research was supported by the Army Research Office under Grant Number W911NF-20-1-0080 and partially funded by ONR Contract N00014-19-1-2620.
+This research was partly sponsored by the U.S.
+National Science Foundation (NSF) under Grant No.
+2416728, Army Research Office under Grant Number W911NF-20-1-0080, and ONR Contract N00014-23-1-2364.
+The views and conclusions contained in this document are those of the authors and should not be interpreted as representing the official policies, either expressed or implied, of the Army Research Office or the U.S.
+Government.
+The U.S.
+Government is authorized to reproduce and distribute reprints for Government purposes notwithstanding any copyright notation herein.
+We thank the authors of Tabsqlify and BINDER for sharing their code, and Yerram Varun for his help in setting up the implementation process.
+We also extend our gratitude to Jennifer Sheffield from UPenn for research support.
 Lastly, we appreciate the reviewing team’s insightful comments.
 Appendix A Supplementary Analysis
-A.1 Comparison With Diverse Methods
-Tables 8 and 9 show results of other baseline methods on WikiTQ and TabFact datasets respectively. Our method, H-STAR, outperformed all other models by substantial margins.
+A.1 FeTaQA
+The results for the FeTaQA dataset using the ROUGE scores Lin (2004) are given in Table 9.
+We observe an incremental improvement in the scores when compared to other methods on the GPT-3.5-Turbo.
+The ROUGE scores focus on lexical similarities while ignoring the semantic similarity between predicted and gold outputs.
+These metrics often fail to capture improvements with in-context learning, as the model doesn’t learn the long-form text style from just an instruction or a few examples.
+Prompting ROUGE-1 ROUGE-2 ROUGE-L
+PaLM-2*
+DATER 0.63 0.41 0.53
+Chain-of-Table 0.66 0.44 0.56
+GPT-3.5-Turbo
+Table-CoT 0.62 0.39 0.51
+TabSQLify 0.58 0.35 0.48
+H-STAR 0.62 0.39 0.52
+Table 9: Performance on FeTaQA dataset. PaLM-2* refers to the Google proprietary PaLM-2 model.
+Figure 7 illustrates an example where the generated output, despite being correct in answering the question, is penalized by the ROUGE metric.
+This highlights the limitations of the metric in evaluating the correctness of the generated responses.
+Figure 7: An example of FeTaQA representing low ROUGE scores for correct answers.
+Qualitative Analysis.
+In this section, we perform a qualitative analysis on 100 samples of FeTaQA with TabSQLify Nahid and Rafiei (2024).
+The study aims to compare the generations of H-STAR with TabSQLify by evaluating the answer quality.
+The answers are evaluated based on their fluency and adequacy compared to the gold answers.
+Table 10 shows that H-STAR has better quality generations in 26% samples, as opposed to only 18% for TabSQLify.
+‘Both’ indicates samples where both the algorithms perform on par.
+This can be attributed to the use of the same LLM.
+Many failures for TabSQLify are derived from an insufficient table, further highlighting the limitations of using a single-view approach for table extraction.
+Method % of samples
+Both 56%
+TabSQLify 18%
+H-STAR 26%
+Table 10: Comparison of H-STAR with TabSQLify on 100 samples.
+Figure 8 presents an example in which TabSQLify returns ‘information not provided’ due to an incorrect table.
+It can be seen that H-STAR generates an answer comparable with the ground truth for the same.
+Figure 9 visually depicts the outputs from both methods.
+It illustrates the evaluation process undertaken to compare the output quality.
+Figure 8: An example of FeTaQA dataset where TabSQLify fails as a result of insufficient table evidence.
+Figure 9: Comparison of the outputs with TabSQLify.
+Human Evaluation.
+We further evaluate the quality of our generations using human evaluation.
+We refer to Nan et al.
+(2022) and evaluate our generations on four factors: (1) Fluency based on the nature and grammar of the output; (2) Correctness measuring the degree of correctness of the prediction; (3) Adequacy to measure whether all the necessary information is contained in the output; (4) Faithfulness to check whether the generation is grounded in the final extracted table used for reasoning.
+We ask five internal human annotators to assign a score of 1-5 for each criterion and report the percentage of samples having values of 4 or 5.
+Table 11 shows the results by five annotators for our method and compares it with other prior methods.
+The performance pattern of H-STAR is similar to human performance with comparable fluency and faithfulness while matching the trend of having higher adequacy over correctness.
+Although we marginally outperform other methods, there is still scope for improvement.
+Method Fluent Correct Adequate Faithful
+T5-large 94.6 54.8 50.4 50.4
+Human 95 92.4 95.6 95.6
+TableCoT 96 82 75 87
+TabSQLify 97 88 84 93
+H-STAR 96.6 87.6 89.6 94
+Table 11: Human evaluation results on FeTaQA dataset.
+A.2 Comparison With Diverse Methods
+Tables 12 and 13 show results of other baseline methods on WikiTQ and TabFact datasets respectively.
+Our method, H-STAR, outperformed all other models by substantial margins.
+Despite using the smaller, less capable GPT-4o-mini, H-STAR outperforms E5 and ReAcTable which use GPT-4 on both TabFact and WikiTQ.
+Furthermore, H-STAR (GPT-3.5-Turbo) outperforms E5 (GPT-4) and ReAct (GPT-4) on WikiTQ.
 Model Method Accuracy
@@
 GPT-3.5-Turbo ReAcTable 52.5
+E5 50.9
 TableCoT 52.4
@@
 SYNTQA(GPT)* 65.2
+GPT-4 ReAcTable 57.3
+E5 65.5
 GPT-3.5-Turbo H-STAR 69.6
+GPT-4o-Mini 74.9
 Table 12: Comparison of H-STAR with additional methods on WikiTQ dataset.
-* SYNTQA (GPT) uses GPT-3.5-Turbo to choose between SQL/text reasoning and uses: 1) fine-tuned models for reasoning; 2) GPT-3.5-Turbo for reasoning.
+* SYNTQA (GPT) uses GPT-3.5-Turbo to choose between SQL/text reasoning and uses: (1) fine-tuned models for reasoning; (2) GPT-3.5-Turbo for reasoning.
 Model Method Accuracy
@@
 GPT-3.5-Turbo ReAcTable 73.1
+E5 75.6
 TableCoT 73.1
 ALTER 84.3
+GPT-4 ReAcTable 83.7
+E5 88.8
 GPT-3.5-Turbo H-STAR 85.0
+GPT-4o-Mini 89.4
 Table 13: Comparison of H-STAR with additional methods on TabFact dataset.
-A.2 FeTaQA
-Represented in Table 10 are the results for the FeTaQA dataset using the ROUGE scores Lin (2004).
-We observe an incremental improvement in the scores when compared to other methods on the GPT-3.5-Turbo.
-The ROUGE scores focus on lexical similarities while ignoring the semantic similarity between predicted and gold outputs.
-These metrics are known to be insensitive to capturing improvements when using in-context learning since the model does not learn the expected style of the long-form text just from an instruction or a few examples.
-Prompting ROUGE-1 ROUGE-2 ROUGE-L
-PaLM-2*
-DATER 0.63 0.41 0.53
-Chain-of-Table 0.66 0.44 0.56
-GPT-3.5-Turbo
-Table-CoT 0.62 0.39 0.51
-TabSQLify 0.58 0.35 0.48
-H-STAR 0.62 0.39 0.52
-Table 10: Performance on FeTaQA dataset. PaLM-2* refers to the Google proprietary PaLM-2 model.
-Figure 7 illustrates an example where the generated output, despite being correct in answering the question, is penalized by the ROUGE metric.
-This highlights the limitations of the metric in evaluating the correctness of the generated responses.
-Figure 7: An example of FeTaQA representing low ROUGE scores for correct answers.
-Human Evaluation
-We further evaluate the quality of our generations using human evaluation.
-We refer to Nan et al.
-(2022) and evaluate our generations on four factors: 1) Fluency based on the nature and grammar of the output; 2) Correctness measuring the degree of correctness of the prediction; 3) Adequacy to measure if all the necessary information is contained in the output; 4) Faithfulness to check whether the generation is grounded to the final extracted table used for reasoning.
-We ask five internal human annotators to assign a score of 1-5 for each criterion and report the percentage of samples having values of 4 or 5.
-Table 11 shows the results by five annotators for our method and compares it with other prior methods.
-The performance pattern of H-STAR is similar to human performance with comparable fluency and faithfulness while matching the trend of having higher adequacy over correctness.
-While we marginally outperform other methods, there is still scope for improvement.
-Method Fluent Correct Adequate Faithful
-T5-large 94.6 54.8 50.4 50.4
-Human 95 92.4 95.6 95.6
-TableCoT 96 82 75 87
-TabSQLify 97 88 84 93
-H-STAR 96.6 87.6 89.6 94
-Table 11: Human evaluation results on FeTaQA dataset.
-Qualitative Analysis
-In this section, we perform a qualitative analysis on 100 samples of FeTaQA with TabSQLify Nahid and Rafiei (2024).
-The study aims to compare the generations of H-STAR with TabSQLify by evaluating the answer quality.
-The answers are evaluated based on their fluency and adequacy when compared with the gold answers.
-Table 12 shows that H-STAR has better quality generations in 26% samples, as opposed to only 18% for TabSQLify.
-‘Both’ indicates samples where both the algorithms perform on par.
-This can be attributed to the use of the same LLM.
-Many failures for TabSQLify are derived from an insufficient table, which further highlights the limitations of using a single-view approach for table extraction.
-Method % of samples
-Both 56%
-TabSQLify 18%
-H-STAR 26%
-Table 12: Comparison of H-STAR with TabSQLify on 100 samples.
-Figure 8 showcases such an example where TabSQLify returns ‘information not provided’ because of incorrect of an extracted table.
-It can be noted that H-STAR generates an answer comparable with the ground truth for the same.
-Figure 9 visually depicts the outputs from both methods.
-It illustrates the evaluation process undertaken to compare the quality of the outputs.
-Figure 8: An example of FeTaQA dataset where TabSQLify fails as a result of insufficient table evidence.
-Figure 9: Comparison of the outputs with TabSQLify.
-A.3 Analysis: Generated Samples
-In Table 13, we study H-STAR by considering the total number of samples generated by the LLMs.
-BINDER and DATER utilize self-consistency techniques to refine their results, while the Chain-of-Table approach adopts a more iterative sample creation process.
-Specifically, BINDER generates 50 Neural-SQL samples, implementing self-consistency, while DATER employs self-consistency at each step, resulting in 100 samples.
-In contrast, the Chain-of-Table method opts for a more resource-efficient strategy, producing 25 samples across three steps: ‘Dynamic Plan’, ‘Generate Args’, and ‘Query’.
-TabSQLify generates the fewest samples, with one generation each for the table decomposition and query steps.
-Using single outputs can result in a decrease in accuracy when the model fails to return an answer.
-Thus, H-STAR incorporates self-consistency measures by generating two outputs at each stage to ensure output validity.
-Each step, except Query, involves two generations for both SQL and Text, i.e., a total of four steps for column and row retrieval.
-These additional steps serve as a precaution against abnormal LLM generations.
-Overall, H-STAR is efficient and only uses 6-10 sample generations.
-Method # samples / step Total # samples
-BINDER Neural SQL: 50 50
-DATER Decompose Table: 40 100
-Generate Cloze: 20
-Generate SQL: 20
-Query: 20
-Chain-of-Table Dynamic Plan $\leq$ 5 $\leq$ 25
-Generate Args $\leq$ 19
-Query: 1
-TabSQLify Table Decompose: 1 2
-Query: 1
-H-STAR Column Extraction: 2-4 6-10
-Row Extraction: 2-4
-Query: 2
-Table 13: Number of generated samples for different methods.
-A.4 Benefit of using Transposed Table
-Table 14 compares H-STAR’s performance using transposed versus original tables for text-based column verification.
-On TabFact, H-STAR achieves 86.51% accuracy with transposed tables, versus 85.22% with original tables.
-On WikiTQ, accuracy is 68.62% with transposed tables, compared to 66.59% with original tables.
+A.3 Advantages of using Transposed Table
+Table 14 compares the performance of H-STAR using original versus transposed tables for text-based column verification.
+In TabFact, H-STAR achieves 86.51% accuracy with transposed tables versus 85.
+22% with the original tables.
+On WikiTQ, accuracy is 68.62% with transposed tables, compared to 66.
+59% with the original tables.
 Method TabFact WikiTQ
@@
 H-STAR (with original table) 85.22 66.59
-Table 14: Comparison between H-STAR using a) transposed table and b) original table for text-based column extraction.
+Table 14: Comparison between H-STAR using (a) transposed table and (b) original table for text-based column extraction.
 Analysis.
 Transposing tables leads to better performance than using the original table for column verification.
-The transposed table approach is more effective for complex datasets like WikiTQ, which are harder to reason upon and contain longer tables.
-Appendix B Input Table Format
+The transposed table approach is more effective for complex datasets such as WikiTQ, which are harder to reason on and contain longer tables.
+Appendix B Comparison with Other Methods
+While H-STAR shares similarities with ReAct and E5 in combining textual reasoning with tool-based methods, there are key differences.
+Like E5, H-STAR employs table extraction to filter irrelevant data and leverages LLMs’ code generation capabilities for table reasoning.
+However, there are significant differences.
+1. Integration of Symbolic and Semantic Reasoning.
+H-STAR combines SQL-based and text-based reasoning, while ReAct merges textual reasoning with actions.
+While ReAct uses tools primarily for actions and E5 applies code generation only for table extraction, both rely solely on semantic approaches for reasoning.
+In contrast, H-STAR applies SQL for quantitative tasks and text for lexical queries.
+2. Fixed Multi-Stage Approach.
+ReAct prompts LLMs to alternate between reasoning and actions, allowing interaction with external sources.
+H-STAR, on the other hand, splits tabular reasoning into two stages: (1) Table Extraction, and (2) Adaptive Reasoning, blending SQL and text reasoning.
+This hybrid approach sets H-STAR apart from ReAct’s focus on textual reasoning and interaction with the environment.
+3. Table Extraction Approach.
+H-STAR distinguishes itself by combining symbolic and semantic methods for table extraction, rather than relying exclusively on either approach.
+Furthermore, H-STAR uses a ‘multi-view’ approach for table extraction, identifying relevant columns from both original and transposed tables before selecting rows, reducing irrelevant data, and improving focus.
+In contrast, ReAct does not employ any extraction method.
+4. Adaptive Reasoning Strategy.
+Unlike E5, H-STAR uses SQL to support semantic reasoning for quantitative tasks while relying solely on semantic reasoning for lexical queries.
+In contrast, ReAct interleaves reasoning and actions without adapting to task type.
+H-STAR’s hybrid approach, combining SQL logic and text understanding, excels at handling diverse tasks, outperforming ReAct on table reasoning tasks.
+Appendix C Input Table Format
 The input table format changes depending on the type of reasoning used.
@@
 (2024).
-B.1 Symbolic Reasoning
+C.1 Textual Reasoning
+We convert the tables into linear, sequential text.
+Continuing with the format of previous textual reasoning methods Ye et al.
+(2023); Chen (2023); Wang et al.
+(2023), we adopt the PIPE encoding i.e.
+plain text separated by ‘|’.
+Furthermore, we append the table with the caption and the list of columns similar to the format in Section C.2.
+Example of the input table:
+table caption: 2012–13 Exeter City F.C. season
+col : name | league | total
+row 1: danny coles | 3 | 3
+row 4: john o’flynn | 11 | 12
+row 8: jamie cureton | 20 | 20
+columns: [’name’, ’league’, ’total’]
+C.2 Symbolic Reasoning
 We adopt the table prompt format from previous SQL-based methods Text-to-SQL Rajkumar et al.
@@
 columns: [’name’, ’league’, ’total’]
-B.2 Textual Reasoning
-We convert the tables into linear, sequential text.
-Continuing with the format of previous textual reasoning methods Ye et al.
-(2023); Chen (2023); Wang et al.
-(2023), we adopt the PIPE encoding i.e.
-plain text separated by ‘|’.
-Furthermore, we append the table with the caption and the list of columns similar to the format in Section B.1.
-Example of the input table:
-table caption: 2012–13 Exeter City F.C. season
-col : name | league | total
-row 1: danny coles | 3 | 3
-row 4: john o’flynn | 11 | 12
-row 8: jamie cureton | 20 | 20
-columns: [’name’, ’league’, ’total’]
-Appendix C Comparison with ReAct
-While our method appears to be similar to ReAct in terms of combining textual reasoning and tool-based methods, there are significant differences between H-STAR and ReAct:
-1. Integration of Symbolic and Semantic Reasoning.
-H-STAR combines SQL-based and text-based reasoning, while ReAct merges textual reasoning with actions.
-ReAct uses tools mainly for actions, not reasoning.
-H-STAR applies SQL for quantitative tasks and text for lexical queries, whereas ReAct focuses more on textual reasoning.
-2. Fixed Multi-Stage Approach.
-ReAct prompts LLMs to alternate between reasoning and actions, allowing interaction with external sources.
-H-STAR, on the other hand, splits tabular reasoning into two stages: 1) Table Extraction, and 2) Adaptive Reasoning, blending SQL and text reasoning.
-This hybrid approach sets H-STAR apart from ReAct’s focus on textual reasoning and interaction with the environment.
-3. Table Extraction Approach.
-H-STAR uses a "multi-view" approach for table extraction, identifying relevant columns from both original and transposed tables before selecting rows, reducing irrelevant data, and improving focus.
-In contrast, ReAct does not employ any extraction method.
-4. Adaptive Reasoning Strategy.
-H-STAR uses SQL to support semantic reasoning for quantitative tasks while relying solely on semantic reasoning for lexical queries.
-In contrast, ReAct interleaves reasoning and actions without adapting to task type.
-H-STAR’s hybrid approach, combining SQL logic and text understanding, excels in handling diverse tasks, outperforming ReAct on table reasoning tasks.
 Appendix D Implementation Details
 We provide the implementation details including the hyperparameters and prompt details for each of the individual steps.
-H-STAR operates in two key stages: 1) Table Extraction and 2) Adaptive Reasoning.
+H-STAR operates in two key stages: (1) Table Extraction and (2) Adaptive Reasoning.
 The Table Extraction phase includes column extraction (SQL-based: colsql and text-based: coltext) and row extraction (SQL-based: rowsql and text-based: rowtext).
 The Adaptive Reasoning phase consists of SQL-based (fsql) and text-based (ftext) reasoning components.
-D.1 Hyperparameters
+D.1 Input Prompts
+Figures 10, 11, 12, 13, 14 and 15 illustrate the input prompts containing sample in-context learning examples for the steps colsql, coltext, rowsql, rowtext, fsql, and ftext respectively.
+D.2 Hyperparameters
 Table 15 provides the details of PaLM-2 hyper-parameters used for the WikiTQ and TabFact datasets.
@@
 Table 16: Hyperparameter settings for GPT-3.5-Turbo.
-D.2 Input Prompts
-Figures 10, 11, 12, 13, 14 and 15 illustrate the input prompts containing sample in-context learning examples for the steps colsql, coltext, rowsql, rowtext, fsql, and ftext respectively.
 Figure 10: Prompt for SQL-based column extraction (colsql).
@@
 Figures 16 and 17 showcase how semantic and symbolic approaches can individually fail for the column and row extraction steps.
-Figures 19, 18, 20 and 21 illustrate error cases for each individual step within H-STAR.
+Figures 18, 19, 20 and 21 illustrate error cases for each individual step within H-STAR.
 Figure 16: Comparison of SQL-based and text-based column extraction methods for answering table queries: (a) SQL-based selection accurately identifies the columns ("1940/41", "description losses"), while text-based selection mistakenly includes "murdered" as a table column.